### v1

# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

Abstract.
We formalized a complete proof of the Auslander–Buchsbaum–Serre criterion in the Lean4 theorem prover. For a local ring, rather than following the well-known proof that considers the residue field as the quotient of the ring by a regular sequence to compute projective dimension and uses the Koszul complex to show the dimension of the cotangent space is at most the global dimension, we prove the criterion via maximal Cohen–Macaulay modules and a weakened version of the Ferrand–Vasconcelos theorem, which is more amenable to the formalization process and the current development of mathlib. Our formalization includes the construction of depth and of Cohen–Macaulay modules and rings, which are used frequently in the proof of the criterion. We also developed related results, including the unmixedness theorem for Cohen–Macaulay rings and Hilbert’s Syzygy theorem.
1. Introduction
The Auslander–Buchsbaum–Serre criterion is a beautiful piece of mathematics which states that for a Noetherian local ring $R$, $R$ is a regular local ring if and only if it has finite global dimension. This criterion did not appear all at once: initially only the “only if” direction was known. At that time, regularity of the localization of a regular local ring at a prime ideal was still an open problem. It was not until 1955 that Jean–Pierre Serre gave the proof of the converse, completing the homological characterization of regular local rings. With this characterization in hand, it is easy to prove that the localization of a regular local ring is again regular. This is one of the earliest applications of homological methods to prove a non-homological result. Furthermore, as a corollary, regularity at all maximal ideals and at all prime ideals coincides, giving a clearer understanding of regular rings.
This project began with the goal of developing the theories of depth and Cohen–Macaulay rings. As the concepts of regular element and regular sequence were already formalized, this goal was within reach given the current library mathlib [7]. The effort was not only of mathematical interest, but also a test of the newly developed homological algebra in mathlib via its application to commutative algebra. As our focus shifted toward combination with homological algebra, we set Hilbert’s Syzygy theorem and the Auslander–Buchsbaum–Serre criterion as targets. However, the $\operatorname{Tor}$ functor was still largely missing from mathlib, and development of the Koszul complex was also incomplete, so the most well-known proof failed on both fronts. The turning point was finding an approach to the global dimension of a regular local ring by proving that a maximal Cohen–Macaulay module over a regular local ring is free; then the target became attainable. Finally, by observing a proof of the Ferrand–Vasconcelos theorem (see [1, Theorem 2.2.8]) showing that induction on dimension can still be performed when focusing only on the unique maximal ideal, we completed the full Auslander–Buchsbaum–Serre criterion. The value of this work is not only in the final theorems: throughout the process we enriched basic constructions in commutative algebra and homological algebra and tested their design and lemmas directly in subsequent developments, making them more user-friendly. We also plan to upstream most of our work into mathlib in the future.
Outline of the Paper
In Section  2 we introduce preliminary results needed for the formalization, including the existing theories in mathlib and the supplementary results we added. We then develop results about depth and Cohen–Macaulay modules and rings in Sections  3 and  4, following [4, Section 6]. This was the earliest part to be formalized. For the theory of regular local rings, Section  5 gives the definition of a regular local ring and basic results about it, such as that a regular local ring is a domain. The Auslander–Buchsbaum–Serre criterion is split into two subsections:  5.2 focuses on proving that the global dimension of a regular local ring equals its Krull dimension, and  5.3 proves a weakened version of the Ferrand–Vasconcelos theorem that implies the converse. These two subsections roughly follow a modified proof obtained from combining [6] and [1, Section 2.2]. In subsection  5.4 we state our final results and their corollaries. In Section  6 we describe implementation issues, related work, and directions for future work.
Throughout the paper, we use the symbol W for external links. Most statements and definitions are accompanied by such a link directly to the source code for the corresponding statement in mathlib or in a branch of our fork W. To keep the links stable, the links to the latter point to a fixed commit of the master branch (shortly after cleanup of the project); other links point to the documentation of mathlib. In some code excerpts the dot notation is exhibited: for example, for I of type Ideal R, the call Ideal.depth I M can be shortened to I.depth M.
Acknowledgements
We would like to thank Professor Liang Xiao for the preliminary review and for organizing the grouped work that made this project possible. We also extend our gratitude to Dr. HuanHuan Yu for producing the blueprint of the formalization project. We are also deeply grateful to all the contributors and reviewers who have been involved in the development and update process of both Lean and mathlib, whose efforts have made our work possible. Special thanks go to Professor Joël Riou for developing the derived category and the $\operatorname{Ext}$ functor, which enabled the homological methods. This work is supported in part by National Key R&D Program of China grant 2024YFA1014000.
2. Preliminaries
Introduction:
In this section, we review some preliminary results that will be used in later proofs.
2.1. Associated Primes
At the start of this project, we only had the definition associatedPrimes W and basic facts about associated primes in mathlibW. Thanks to the work of Jinzhao Pan, we have the finiteness of associated primes for a finitely generated module over a Noetherian ring in mathlibW.
theorem associatedPrimes.finite (A : Type u) [CommRing A] (M : Type v) [AddCommGroup M] [Module A M] [IsNoetherianRing A] [Module.Finite A M] : (associatedPrimes A M).Finite
Combined with the fact that the complement of the non-zero-divisors of a module is exactly the union of its associated primesW, this lemma lets us apply prime-avoidance arguments when looking for a regular element in an ideal.
We also formalized results on associated primes of localized modulesW. For $S$ a multiplicative set of a Noetherian ring $R$ and $M$ a finitely generated $R$-module, one has $\mathrm{Ass}(S^{-1}M)=\mathrm{Ass}(M)\cap\operatorname{Spec}(S^{-1}R)$. Denoting $R^{\prime}$ and $M^{\prime}$ the localizations of $R$ and $M$ at $S$ respectively, the main statement is:
lemma preimage_comap_associatedPrimes_eq_associatedPrimes_of_isLocalizedModule [IsNoetherianRing R] :
(Ideal.comap (algebraMap R R’)) ⁻¹’ (associatedPrimes R M) = associatedPrimes R M
Notice that the induced map $\operatorname{Spec}(R^{\prime})\to\operatorname{Spec}(R)$ is injective, this lemma already states the strongest result.
Furthermore, applying the above to the localization at $p$ a minimal prime over $\mathrm{Ann}(M)$: since associated primes of $M_{p}$ exist, any such associated prime must be the maximal ideal of $R_{p}$. From the previous result we deduce that minimal primes over $\mathrm{Ann}(M)$ lies in $\mathrm{Ass}(M)$W.
lemma minimalPrimes_annihilator_subset_associatedPrimes [IsNoetherianRing R] [Module.Finite R M] :
(Module.annihilator R M).minimalPrimes associatedPrimes R M
2.2. Krull Dimension of Module
We define the Krull dimension of a module simply as the Krull dimension of its support viewed as a preorder setW. Some basic results are formalized hereW.
Similar to dimension of ring, for $R$-module $M$ quotienting by an element in the jacobson radical of $\mathrm{Ann}(M)$ reduces the dimension by at most oneW.
theorem supportDim_le_supportDim_quotSMulTop_succ_of_mem_jacobson {x : R} (h : x (annihilator R M).jacobson) :
supportDim R M supportDim R (QuotSMulTop x M) + 1
The formalization of this lemma follows [3, Proposition 2.3.4] via [3, Lemma 2.3.3]W.
If we quotient by an element $x$ that is not contained in any minimal prime over $\mathrm{Ann}(M)$, the dimension decreases by at least oneW:
theorem supportDim_quotSMulTop_succ_le_of_notMem_minimalPrimes {x : R} (hn : p (annihilator R M).minimalPrimes, x p) :
supportDim R (QuotSMulTop x M) + 1 supportDim R M
Combining this with the fact that an $M$-regular element is not contained in any minimal prime over $\mathrm{Ann}(M)$, we obtainW:
theorem supportDim_quotSMulTop_succ_eq_supportDim_mem_jacobson {x : R} (reg : IsSMulRegular M x)
(hx : x (annihilator R M).jacobson) : supportDim R (QuotSMulTop x M) + 1 = supportDim R M
If $R$ is local, induction on the length of a regular sequence shows that quotienting by a regular sequence decreases the dimension of a module by exactly its lengthW:
theorem supportDim_add_length_eq_supportDim_of_isRegular (rs : List R) (reg : IsRegular M rs) :
supportDim R (M ofList rs (Top.top : Submodule R M)) + rs.length = supportDim R M
2.3. Dimension of Polynomial Rings
As we need to work with polynomials over Cohen–Macaulay and regular rings, some results about the heights of ideals in polynomial rings over Noetherian rings are required. Let $R$ be a Noetherian ring, we establish the following results, which follow easily from Krull’s height theoremW.
lemma height_map_maximalIdeal [IsLocalRing R] : ((maximalIdeal R).map C).height = (maximalIdeal R).height
lemma height_of_comap_eq_maximalIdeal [IsLocalRing R] (m : Ideal R[X]) [h : m.IsPrime]
(eqm : m.comap C = maximalIdeal R) : m.height (maximalIdeal R).height + 1
lemma height_le_height_comap_succ (p : Ideal R[X]) [p.IsPrime] : p.height (p.comap C).height + 1
From these we obtain the standard result on the Krull dimension of polynomial ring over a Noetherian ringW.
lemma ringKrullDim_of_isNoetherianRing [IsNoetherianRing R] : ringKrullDim R[X] = ringKrullDim R + 1
There is an ongoing effort to formalize similar statements more generally by studying heights under ring homomorphisms with the going-down property.(SeeWW.) Because that work is more general and these lemmas here are proven specifically for polynomial rings (with some proofs “forced through”), they are intended as temporary and will be deprecated once the more general versions are available in mathlib. See  6.2 for details.
2.4. Regular Sequence Under Flat Base Change
As we need to handle preservation of regular sequences under localization, we formalized the more general results that weakly regular sequences are preserved under flat base changeW and regular sequences are preserved under faithfully flat base changeW.
theorem RingTheory.Sequence.IsRegular.of_faithfullyFlat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S] {M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N] [IsScalarTower R S N] [Module.Flat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R} (reg : IsWeaklyRegular M rs) :
IsWeaklyRegular N (List.map (algebraMap R S) rs)
theorem RingTheory.Sequence.IsRegular.of_faithfullyFlat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S] {M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N] [IsScalarTower R S N] [Module.FaithfullyFlat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R} (reg : IsRegular M rs) :
IsRegular N (List.map (algebraMap R S) rs)
Specializing to the case where the module is the ring itself yields a more user-friendly versionWW.
Finally, we also added specialized lemmas for localized modules and localization, which show that a regular sequence contained in the prime being localized at remains regular after localizationWW.
2.5. Projective Dimension
As the $\operatorname{Ext}$ functor is a relatively recent addition to the development of mathlib (see [5]), many related constructions are not yet fully available. We only have a predicate describing that an object $X$ of an abelian category $C$ has projective dimension less than some $n\in\mathbb{N}$, defined as the vanishing of all $\operatorname{Ext}_{i}(X,Y),\forall Y,\forall n\leq i$
class HasProjectiveDimensionLT (X : C) (n : ℕ) : Prop where mk ::
subsingleton (i : ℕ) (hi : n i) Y : C :
Subsingleton (Ext.{max u v} X Y i)
With only this definition, some operations are awkward (for example adding a number). Even if we try to pick out the minimal $n$ such that HasProjectiveDimensionLT X (n + 1), the disadvantages are significant: first such number may doesn’t exist at all (if $X$ has no finite projective resolution); second, finding it with Nat.find complicates operations and requires decidable instances. For these reasons we want projective dimension as a concrete numeric object.
However, projective dimension cannot be a plain nat: first, we need to give objects with no finite projective resolution value $\infty$, second, by common convention the projective dimension of the zero object is -$\infty$. We therefore define projective dimension in WithBot ℕ∞:
variable {C : Type u} [Category.{v, u} C] [Abelian C]
noncomputable def projectiveDimension (X : C) : WithBot ℕ∞ :=
sInf {n : WithBot ℕ∞ | (i : ℕ), n < i HasProjectiveDimensionLT X i}
Although this definition looks a little verbose, it behaves as in the usual conventions. (Out side code excerpts we denote projectiveDimension by $\operatorname{projdim}$)
With this definition we can prove the equivalence projectievDimension X < n HasProjectiveDimensionLT X n and many other useful lemmas to make the numeric version convenient in practice. Many lemmas related to HasProjectiveDimensionLT now have corresponding statements for projectievDimension.
We also prove a lemma specific to the category of $R$-modulesW.
Lemma 2.1.
For a finitely generated module $M$ over Noetherian ring $R$, $\operatorname{projdim}(M)$ is equal to the supremum of $\operatorname{projdim}(M_{\mathfrak{p}})$ over the prime ideals $\mathfrak{p}$ of $R$.
lemma projectiveDimension_eq_iSup_localizedModule_prime [Small.{v} R] [IsNoetherianRing R]
(M : ModuleCat.{v} R) [Module.Finite R M] :
projectiveDimension M = iSup (p : PrimeSpectrum R), projectiveDimension (M.localizedModule p.1.primeCompl)
For convenience in formalization, we do not directly prove that $\operatorname{Ext}$ commutes with localization. Noting that the left-hand side being -$\infty$ is equivalent to $M=0$ and similarly for the right-hand side -$\infty$, we instead prove that for each natural number $n$, the statement left hand side no greater than $n$ is equivalent to the right hand side no greater than $n$, by induction on $n$. Then we construct an exact sequence $0\rightarrow N\rightarrow F\rightarrow M\rightarrow 0$ with $F$ free, localizing at a prime $\mathfrak{p}$ yields an exact sequence $0\rightarrow N_{\mathfrak{p}}\rightarrow F_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}\rightarrow 0$. Using the existing lemma about projective dimension in short exact sequencesW, we reduce the case $n+1$ to the case $n$. We are then left to show that $M$ is projective if and only if each localization at prime $M_{\mathfrak{p}}$ is projective, which is already in mathlib WW. There is also a version of this lemma for running only over all maximal ideals W.
Similarly we prove the following lemmaW.
Lemma 2.2.
For $S$ a multiplicative set of $R$ and $R$-module $M$, proective dimension of $S^{-1}M$ over $S^{-1}R$ is no greater than projective dimension of $M$ over $R$.
lemma projectiveDimension_le_projectiveDimension_of_isLocalizedModule [Small.{v} R]
(S : Submonoid R) (M : ModuleCat.{v} R) :
projectiveDimension (M.localizedModule S) projectiveDimension M
2.6. Global Dimension
With projective dimension as a numeric object, we can directly define global dimension of a ring $R$ as the supremum of the projectve dimensions over all $R$-modulesW.
noncomputable def globalDimension : WithBot ℕ∞ :=
iSup (M : ModuleCat.{v} R), projectiveDimension.{v} M
Basic properties and the relation with HasProjectiveDimensionLE are formalized in the same fileW.(Outside code excerpts we denote globalDimension by $\operatorname{gldim}$)
The supremum over all modules can actually be reduced to the supremum over all finitely generated modules. However, since in mathlib projective dimension is defined via vanishing of Ext and a definition of injective dimension was missing when this formalization started, we formalized only the following weakened form of [4, Lemma 18.2]W.
Lemma 2.3.
For a commutative ring $R$ and $n$ a non-negative integer the following are equivalent:
  1. (1)
    $\operatorname{gldim}(R)\leq n$
  2. (2)
    $\operatorname{projdim}(M)\leq n$ for all finitely generated $R$-module $M$
  3. (3)
    $\operatorname{Ext}_{R}^{i}(M,N)$ vanishes for all $R$-module $M,N$ for index $i>n$
$(1)\rightarrow(2)$ and $(2)\rightarrow(3)$ are trivial, For $(3)\rightarrow(1)$, the following lemma sufficesW:
Lemma 2.4.
For $R$-module $M$ and $n$ a non-negative integer, if $\operatorname{Ext}_{R}^{n+1}(R/I,M)$ vanish for all $I$ ideal of $R$, then for any $R$-module NN, $\operatorname{Ext}_{R}^{n+1}(N,M)$ vanishes.
lemma ext_subsingleton_of_quotients [Small.{v} R] (M : ModuleCat.{v} R) (n : ℕ)
(h : I : Ideal R, Subsingleton (Ext (ModuleCat.of R (Shrink.{v} (R I))) M (n + 1))) :
N : ModuleCat.{v} R, Subsingleton (Ext N M (n + 1))
This result is essentially about the $n$-th syzygy of $M$ in an injective resolution. We prove it by induction on $n$:using dimension-shifting on $M$ we reduce to the case $n=0$, which is just an application of the existing Baer criterion. W.
Using Lemma  2.3, we show that the global dimension is the supremum of the projective dimensions of finitely generated modulesW.
lemma globalDimension_eq_sup_projectiveDimension_finite [Small.{v, u} R] : globalDimension.{v} R =
iSup (M : ModuleCat.{v} R), iSup (_ : Module.Finite R M), projectiveDimension.{v} M
Combining this with the localization results from the previous subsection, we formalize the following statementW.
Lemma 2.5.
For noetherian ring $R$, $\operatorname{gldim}(R)$ is equal to the supremum of $\operatorname{gldim}(R_{\mathfrak{p}})$ over the prime ideals $\mathfrak{p}$ of $R$.
lemma globalDimension_eq_iSup_loclization_prime [Small.{v} R] [IsNoetherianRing R] :
globalDimension.{v} R = iSup (p : PrimeSpectrum R), globalDimension.{v} (Localization.AtPrime p.1)
Since the left-hand side equals the supremum of projective dimensions of finitely generated modules, Lemma  2.1, we know that the left is no greater than the right. Conversely, any $R_{\mathfrak{p}}$-module $N$ is already $R$-module with $N=N_{\mathfrak{p}}$, and Lemma  2.2 yields the reverse inequality, finishing the proof. A version for localizing only at maximal ideals is also providedW.
3. Definition of Depth and Related Results
Introduction:
In this section we formalize the definition of depth as the minimal index of non-vanishing $\operatorname{Ext}$. Before doing so, and to relate depth to the maximal length of a regular sequence contained in a given ideal, we formalized Rees theorem (see Theorem  3.1). We then establish basic properties of depth, such as the inequalities associated to a short exact sequence and the fact that quotienting by a regular element decreases depth by exactly one. Finally, we present our formalizations of two classical results about depth: the Auslander–Buchsbaum theorem (see 3.2) and Ischebeck’s theorem (see 3.3).
Throughout this section we let $R$ be a Noetherian ring and $M$, NN be $R$-modules.
Theorem 3.1.
(Rees) For $M$ a finitely generated $R$-module, and let $I$ an ideal of $A$ with $IM\neq M$. Fix an integer $n>0$. The following are equivalent.
  1. (1)
    $\operatorname{Ext}^{i}_{R}(N,M)=0$ for all $i<n$, and every finitely generated $R$-module NN with $\mathrm{Supp}(N)\subseteq V(I)$;
  2. (2)
    $\operatorname{Ext}_{R}^{i}(R/I,M)=0$ for all $i<n$;
  3. (3)
    there exists a finite $R$-module NN with $\mathrm{Supp}(N)=V(I)$ such that $\operatorname{Ext}^{i}_{R}(N,M)=0$ for every $i<n$;
  4. (4)
    there exists an $M$-regular sequence $x_{1},\dots,x_{n}$ of length $n$ in $I$.
[4, Theorem 28]
(The equivalence $(2)\longleftrightarrow(4)$ will be a key ingredient in the definition of depth.)
We first prove the following lemma: for finitely generated $M$ and $N$, the $R$-linear maps $N\rightarrow M$ are all zero if and only if there exists a $M$-regular element in the annihilator of MMW.
lemma subsingleton_linearMap_iff [IsNoetherianRing R] [Module.Finite R M] [Module.Finite R N] :
Subsingleton (N →[R] M) r Module.annihilator R N, IsSMulRegular M r
A sketch of the proof : if a prime ideal $\mathfrak{p}\in\mathrm{Ass}(M)$ contains $\mathrm{Ann}(N)$, then the maximal ideal $\mathfrak{m}_{\mathfrak{p}}$ of $R_{\mathfrak{p}}$ is an associated prime of $M_{\mathfrak{p}}$, we can obtain a non-trivial linear map $N_{\mathfrak{p}}\rightarrow N_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}N_{\mathfrak{p}}\rightarrow R_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}$ which yields a nontrivial map $M\rightarrow N$ non-trivial, then applying prime avoidance to $\mathfrak{p}\in\mathrm{Ass}(M)$ finishes the proof.
We now present the proof of the main theorem formalized as followsW.
lemma exist_isRegular_tfae [IsNoetherianRing R] (I : Ideal R) [Small.{v} (R I)] (n : ℕ)
(M : ModuleCat.{v} R) (Mntr : Nontrivial M) (Mfin : Module.Finite R M) (smul_lt : I · (Top.top : Submodule R M) < Top.top) :
[∀ N : ModuleCat.{v} R, (Nontrivial N Module.Finite R N
Module.support R N PrimeSpectrum.zeroLocus I) i < n, Subsingleton (Ext N M i),
i < n, Subsingleton (Ext (ModuleCat.of R (Shrink.{v} (R I))) M i),
N : ModuleCat R, Nontrivial N Module.Finite R N
Module.support R N = PrimeSpectrum.zeroLocus I i < n, Subsingleton (Ext N M i),
rs : List R, rs.length = n (∀ r rs, r I) RingTheory.Sequence.IsRegular M rs
].TFAE
The implications $(1)\rightarrow(2)$ and $(2)\rightarrow(3)$ are straightforward. The implications $(3)\rightarrow(4)$ and $(4)\rightarrow(1)$ follow the arguments in [4, Theorem 28]. The lemma above plays an important role in both directions and is therefore separated as an independent result.
With Rees’ theorem established, we proceed to the definition of depth.
3.1. Definition of Depth
The Rees theorem above builds a bridge between the homological characterization and the existence of regular sequences. We would like a definition of $\operatorname{depth}$ that is meaningful both in commutative algebra and homological terms. We therefore adopt the following definition (often called the ‘grade’ in some texts)W.
def moduleDepth (N M : ModuleCat.{v} R) : ℕ∞ :=
sSup {n : ℕ∞ | i : ℕ, i < n Subsingleton (Ext N M i)}
Here $R$ is a commutative ring of Type u that is Small.{v} to ensure ModuleCat.{v} R has enough projectives. These assumptions on $R$ are used throughout in this section. (out side code excerpts, moduleDepth M N will be denoted by $\operatorname{depth}(M,N)$)
This definition may look unusual but matches the intended idea: it is the minimal index of non-vanishing $\operatorname{Ext}_{R}^{i}(N,M)$, and it equals $\infty$ when no such index exists.
We now specialize to the cases of primary interest. The $\operatorname{depth}$ of a module with respect to an ideal is defined as the followingW:
def Ideal.depth (I : Ideal R) (M : ModuleCat.{v} R) : ℕ∞ :=
moduleDepth (ModuleCat.of R (Shrink.{v} (R I))) M
For a local ring, the $\operatorname{depth}$ of a module with respect to the unique maximal ideal isW:
def IsLocalRing.depth [IsLocalRing R] (M : ModuleCat.{v} R) : ℕ∞ :=
(IsLocalRing.maximalIdeal R).depth M
Using Rees theorem, we relate the above homological notion to the maximal length of $M$-regular sequence contained in a given ideal. In the statements below we assume $M$ and $N$ are nontrivial, finitely generated $R$-modules.
If $I$ an ideal of $R$ with $IM<M$ and $\mathrm{Supp}(N)=V(I)$, then $\operatorname{depth}(N,M)=\operatorname{depth}_{I}(M)$W.
lemma moduleDepth_eq_depth_of_supp_eq [IsNoetherianRing R] (I : Ideal R)
(N M : ModuleCat.{v} R) [Module.Finite R M] [Nfin : Module.Finite R N]
[Nontrivial M] [Nntr : Nontrivial N] (smul_lt : I · (Top.top : Submodule R M) < Top.top)
(hsupp : Module.support R N = PrimeSpectrum.zeroLocus I) :
moduleDepth N M = I.depth M
Under the same hypotheses, $\operatorname{depth}(N,M)$ equals the supremum of lengths of $M$-regular sequences contained in IIW:
lemma moduleDepth_eq_sSup_length_regular [IsNoetherianRing R] (I : Ideal R)
(N M : ModuleCat.{v} R) [Module.Finite R M] [Nfin : Module.Finite R N]
[Nontrivial M] [Nntr : Nontrivial N] (smul_lt : I · (Top.top : Submodule R M) < Top.top)
(hsupp : Module.support R N = PrimeSpectrum.zeroLocus I) :
moduleDepth N M = sSup {(List.length rs : ℕ∞) | (rs : List R)
(_ : RingTheory.Sequence.IsRegular M rs) (_ : r rs, r I) }
With the homological definition in hand, we obtain the following inequalities for a short exact sequence of nontrivial finitely generated $R$-modules.
Theorem 3.2.
For a short exact sequence of non-trivial finitely generated $R$ modules $0\rightarrow M^{\prime}\rightarrow M\rightarrow M^{\prime\prime}\rightarrow 0$
  1. (1)
    $\operatorname{depth}(N,M)\geq min\{\operatorname{depth}(N,M^{\prime}),\operatorname{depth}(N,M^{\prime\prime})\}$
  2. (2)
    $\operatorname{depth}(N,M^{\prime\prime})\geq min\{\operatorname{depth}(N,M),\operatorname{depth}(N,M^{\prime})-1\}$
  3. (3)
    $\operatorname{depth}(N,M^{\prime})\geq min\{\operatorname{depth}(N,M),\operatorname{depth}(N,M^{\prime\prime})+1\}$
A generalization of [6, 00LX]
These follow readily from the long exact sequence and can be found hereW.
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(N,M^{\prime})\rightarrow\operatorname{Ext}_{R}^{i}(N,M)\rightarrow\operatorname{Ext}_{R}^{i}(N,M^{\prime\prime})\rightarrow\operatorname{Ext}_{R}^{i+1}(N,M^{\prime})\rightarrow\operatorname{Ext}_{R}^{i+1}(N,M)\rightarrow\cdots$
The corresponding dual statements about $\operatorname{depth}(\cdot,M)$ in short exact sequence in short exact sequences are proved similarly using the contravariant long exact sequence, seeW. However for Ideal.depth and IsLocalRing.depth the dual version is rarely used.
We also formalized the behavior of depth under quotient by a regular element.
Theorem $3.$
For a $M$-regular element $x\in\mathrm{Ann}(N)$, we have $\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$
A generalization of [6, 090R]
Indeed, from the short exact sequence $0\rightarrow M\xrightarrow{\cdot x}M\rightarrow M/xM\rightarrow 0$, we obtain long exact sequence
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i}(N,M)\rightarrow\operatorname{Ext}_{R}^{i}(N,M/xM)\rightarrow\operatorname{Ext}_{R}^{i+1}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i+1}(N,M)\rightarrow\cdots$
Since $x\in Ann(N)$, the maps $\operatorname{Ext}_{R}^{i}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i}(N,M)$ are zero for all $i$, so $\operatorname{Ext}_{R}^{i}(N,M/xM)$ vanish if and only if both $\operatorname{Ext}_{R}^{i}(N,M)$ and $\operatorname{Ext}_{R}^{i+1}(N,M)$ vanish. From this the theorem followsW.
lemma moduleDepth_quotSMulTop_succ_eq_moduleDepth (N M : ModuleCat.{v} R) (x : R)
(reg : IsSMulRegular M x) (mem : x Module.annihilator R N) :
moduleDepth N (ModuleCat.of R (QuotSMulTop x M)) + 1 = moduleDepth N M
Here QuotSMulTop x M denotes the quotient module $M/xM$.
We also formalized the analogous statement for quotient by a (weakly) regular sequence. This follows by induction on the length of the sequenceW.
lemma moduleDepth_quotient_regular_sequence_add_length_eq_moduleDepth (N M : ModuleCat.{v} R)
(rs : List R) (reg : IsWeaklyRegular M rs) (h : r rs, r Module.annihilator R N) :
moduleDepth N (ModuleCat.of R (M (Ideal.ofList rs) · (Top.top : Submodule R M))) + rs.length = moduleDepth N M
3.2. The Auslander–Buchsbaum Theorem
In this section we prove the following theorem.
Theorem 3.4.
(Auslander–Buchsbaum) Let $(R,\mathfrak{m},k)$ be a Noetherian local ring and let $M$ be a non-trivial finitely generated $R$-module. If projective dimension of $M$ is finite, then
$\operatorname{projdim}(M)+\operatorname{depth}(M)=\operatorname{depth}(A)$
We first establish a preliminary result. Since we formalized that $\operatorname{Ext}$ commute with finite bi-productW, for a finitely generated free module $M$ over $R$, $\operatorname{Ext}_{R}^{i}(N,M)$ vanish if and only if $\operatorname{Ext}_{R}^{i}(N,R)$ vanish for any $R$-module NNW. As a corollary $\operatorname{depth}(M)=\operatorname{depth}(R)$W.
lemma finte_free_ext_vanish_iff
(M N : ModuleCat.{v} R) [Module.Finite R M] [Module.Free R M] [Nontrivial M] (i : ℕ) :
Subsingleton (Ext.{max u v} N M i) Subsingleton (Ext.{max u v} N (ModuleCat.of R (Shrink.{v} R)) i)
lemma free_depth_eq_ring_depth (M N : ModuleCat.{v} R) [Module.Finite R M] [Module.Free R M] [Nontrivial M] :
moduleDepth N M = moduleDepth N (ModuleCat.of R (Shrink.{v} R))
Back to the main theorem, the case $\operatorname{projdim}(M)=0$ is immediate since then $M$ is projective and hence free.
The case $\operatorname{projdim}(M)=1$ is the core, so we prove it separately.
lemma AuslanderBuchsbaum_one [IsNoetherianRing R] [IsLocalRing R] (M : ModuleCat.{v} R) [Nontrivial M]
[Module.Finite R M] (le1 : HasProjectiveDimensionLE M 1) (nle0 : ¬ HasProjectiveDimensionLE M 0) :
1 + IsLocalRing.depth M = IsLocalRing.depth.{v} (ModuleCat.of.{v} R (Shrink.{v} R))
Pick a $k$ basis of $M/\mathfrak{m}M$ and lift it back to $M$, so we obtain a map $f:R^{n}\rightarrow M$ whose induced map $(R/\mathfrak{m})^{n}\rightarrow M/\mathfrak{m}M$ is isomorphism, hence $\mathrm{ker}(f)\subseteq\mathfrak{m}R^{n}$. Because $\mathrm{ker}(f)$ has projective dimension 0, it is projective thus free. By the following lemma, we know that the inclusion from $\mathrm{ker}(f)$ to $R^{n}$ is contained in $\mathfrak{m}(\mathrm{ker}(f)\rightarrow R^{n})$ (view $R$-linear maps as $R$ module).
theorem mem_smul_top_of_range_le_smul_top [Module.Finite R M] [Module.Free R M] (f : M →[R] N)
(hf : LinearMap.range f I · (Top.top : Submodule R N)) : f I · (Top.top : Submodule R (M →[R] N))
Applying the next lemma with $L=R/\mathfrak{m}$, the map $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})$ induced by the inclusion $\mathrm{ker}(f)\rightarrow R^{n}$ is zero.
lemma ext_hom_zero_of_mem_ideal_smul (L M N : ModuleCat.{v} R) (n : ℕ) (f : M N)
(mem : f (Module.annihilator R L) · (Top.top : Submodule R (M N))) :
(AddCommGrp.ofHom <| ((Ext.mk f)).postcomp L (add_zero n)) = 0
(One easily verifies this lemma by reducing to the case $f=a\cdot g$ with $a\in\mathrm{Ann}(L)$.)
Now consider the long exact sequence induced by $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$.
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})\rightarrow\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i+1}(k,R^{n})\rightarrow\cdots$
The first and fourth maps are zero by the discussion above. Moreover, vanishing of $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))$ and of $\operatorname{Ext}_{R}^{i}(k,R^{n})$ are each equivalent to vanishing of $\operatorname{Ext}_{R}^{i}(k,R)$. From the above long exact sequence we deduce that $\operatorname{Ext}_{R}^{i}(k,M)$ vanishes if and only if both $\operatorname{Ext}_{R}^{i}(k,R)$ and $\operatorname{Ext}_{R}^{i+1}(k,R)$ vanish. By the homological characterization of depth this yields $1+\operatorname{depth}(M)=\operatorname{depth}(A)$.
Finally, we prove the theorem by induction on $\operatorname{projdim}(M)$. Assume the statement holds for projective dimension $n\geq 1$, and let $\operatorname{projdim}(M)=n+1$. Take a surjection $f:R^{m}\rightarrow M$, then $\mathrm{ker}(f)$ has projective dimension $n$. By induction hypothesis, $n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$. Since $\operatorname{depth}(R^{m})=\operatorname{depth}(R)\geq n>0$, so there is no non-trivial $k\rightarrow\mathrm{ker}(f)\rightarrow R^{m}$, thus $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$ vanish.
Consider the long exact sequence associated to $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$.
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})\rightarrow\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i+1}(k,R^{n})\rightarrow\cdots$
For all $i<\operatorname{depth}(\mathrm{ker}(f))<\operatorname{depth}(R)$, the map $\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))$ is an isomorphism. Together with $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$ vanish, we obtain $\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(M)+1$. Hence $n+1+\operatorname{depth}(M)=n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$, which completes the induction and the proof.
3.3. The Ischebeck Theorem
With the general definition moduleDepth, we can state the Ischebeck theorem as follows: for Noetherian local ring $R$ and non-trivial finitely generated module $M$ and $N$, $\operatorname{depth}(M,N)\geq\operatorname{depth}(M)-\dim(N)$W.
theorem moduleDepth_ge_depth_sub_dim [IsNoetherianRing R] [IsLocalRing R] (M N : ModuleCat.{v} R)
[Module.Finite R M] [Nfin : Module.Finite R N] [Nontrivial M] [Nntr : Nontrivial N]
[Small.{v} R] : moduleDepth N M IsLocalRing.depth M -
(Module.supportDim R N).unbot (Module.supportDim_ne_bot_of_nontrivial R N)
Since the case $\dim(N)=\infty$ is automatically correct, we prove the statement by strong induction on $\dim(N)$. If $\dim(N)=0$, then $\mathrm{Supp}(N)=V(\mathfrak{m})$, and hence $\operatorname{depth}(N,M)=\operatorname{depth}(M)$ by results established earlier in this section.
Now assume the statement holds for modules of dimension no greater than $n$, and let $\dim(N)=n+1$. In the development of associated primes one has an induction principle for finitely generated modules over a Noetherian ringW. It suffices to check the property in the following cases: $N=0$ (irrelevant here), $N\cong R/\mathfrak{p}$ for some prime ideal $\mathfrak{p}$, and that the property is preserved under short exact sequences. Preservation under exact sequences follows easily from Theorem 3.2 and a straightforward calculation. Be careful to apply the induction hypothesis appropriately to $N_{1}$ in the two subcases when $\dim(N_{1})\leq n$ and when $\dim(N_{1})=n+1$, because in this process the cited induction principle only provides hypotheses for modules of dimension $n+1$. The hypotheses in the former case come from the strong induction on dimension (and similarly for $N_{3}$). Thus the remaining and essential case is $N\cong R/\mathfrak{p}$, which is handled by the same argument as in [4, Lemma 15.2].
Using this theorem, we obtain a useful bound on $\operatorname{depth}$ by taking $N=R/\mathfrak{p}$ for $\mathfrak{p}\in\mathrm{Ass}(M)$. Since $\operatorname{depth}(M,R/\mathfrak{p})=0$, we have $\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})$W. See also [4, Theorem 29].
theorem depth_le_ringKrullDim_associatedPrime [IsNoetherianRing R] [IsLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] [Nontrivial M] (P : Ideal R) (ass : P associatedPrimes R M) :
IsLocalRing.depth M (ringKrullDim (R P)).unbot (quotient_prime_ringKrullDim_ne_bot ass.1)
Moreover, from this we deduce the familiar inequalities $\operatorname{depth}(M)\leq\dim(M)\leq\dim(R)<\inftyW$.
4. Cohen–Macaulay Modules and Rings
Introduction:
In this section, we first give the definition of a Cohen–Macaulay module over a local ring. Then we prove that passing to the quotient by a regular element preserves the Cohen–Macaulay property, and that localization of a Cohen–Macaulay module is still Cohen–Macaulay. Similarly, we can establish the corresponding results for Cohen–Macaulay rings. In  4.1 we prove the unmixed theorem for Cohen–Macaulay rings; as a by‑product we obtain that a Cohen–Macaulay local ring is catenary. Since a polynomial ring over a Cohen–Macaulay ring is still Cohen–Macaulay (proved in  4.2), it will be straightforward to deduce that a Cohen–Macaulay ring is universally catenary once the preliminary constructions are in place.
With a well-developed definition of depth, we give the definition of a Cohen–Macaulay moduleW (over local ring) following [4, 16.A]. (In this section, let $R$ be a commutative ring, not assuming it to be Noetherian at first.)
class ModuleCat.IsCohenMacaulay [IsLocalRing R] [Small.{v} R] (M : ModuleCat.{v} R) : Prop where
depth_eq_dim : Subsingleton M Module.supportDim R M = IsLocalRing.depth M
For a finitely generated Cohen–Macaulay module $M$ over a Noetherian local ring $R$, for any $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $\mathrm{Ann}(M)\subseteq\mathfrak{p}$. As shown in  3.3, we know $\dim(M)=\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})\leq\dim(R/\mathrm{Ann}(M))=\dim(M)$, so all inequalities are equalities. Hence $\dim(M)=\operatorname{depth}(M)=\dim(R/\mathfrak{p})$, forcing $\mathfrak{p}$ to be minimal prime over $\mathrm{Ann}(M)$. This yields the following two lemmasW W.
lemma depth_eq_dim_quotient_associated_prime_of_isCohenMacaulay (M : ModuleCat.{v} R)
[M.IsCohenMacaulay] [Module.Finite R M] [Nontrivial M] (p : Ideal R) (mem : p associatedPrimes R M) :
IsLocalRing.depth M = (ringKrullDim (R p)).unbot (quotient_prime_ringKrullDim_ne_bot mem.1)
lemma associated_prime_minimal_of_isCohenMacaulay (M : ModuleCat.{v} R)
[M.IsCohenMacaulay] [Module.Finite R M] [Nontrivial M] (mem : p associatedPrimes R M) :
p (Module.annihilator R M).minimalPrimes
Combining with the final result in  2.1, we know that the associated primes of a Cohen–Macaulay module are exactly the minimal primes over $\mathrm{Ann}(M)$W.
lemma associated_prime_eq_minimalPrimes_isCohenMacaulay
(M : ModuleCat.{v} R) [M.IsCohenMacaulay] [Module.Finite R M] [Nontrivial M] :
associatedPrimes R M = (Module.annihilator R M).minimalPrimes
From the results mentioned in  3.1 and  2.2, we know that for a finitely generated $R$-module $M$ with $R$ local, quotienting by an $M$-regular element reduces both the depth and the Krull dimension of $M$ by exactly one; therefore whether $M$ is Cohen–Macaulay is preserved under such a quotientW. A similar result holds for quotienting by a regular sequenceW. See also [4, Theorem 30(ii)]
lemma quotSMulTop_isCohenMacaulay_iff_isCohenMacaulay
(M : ModuleCat.{v} R) [Module.Finite R M] (r : R) (reg : IsSMulRegular M r) (mem : r maximalIdeal R) :
M.IsCohenMacaulay (ModuleCat.of R (QuotSMulTop r M)).IsCohenMacaulay
lemma quotient_regular_isCohenMacaulay_iff_isCohenMacaulay
(M : ModuleCat.{v} R) [Module.Finite R M] (rs : List R) (reg : IsRegular M rs) :
M.IsCohenMacaulay (ModuleCat.of R (M Ideal.ofList rs · (Top.top : Submodule R M))).IsCohenMacaulay
We then formalized that localization of a Cohen–Macaulay module at prime ideal is still Cohen–MacaulayW.
lemma isLocalize_at_prime_depth_eq_of_isCohenMacaulay (p : Ideal R) [p.IsPrime] [IsLocalRing Rₚ]
[Module.Finite R M] [M.IsCohenMacaulay] : Mₚ.IsCohenMacaulay
Here $R_{\mathfrak{p}}$ and $M_{\mathfrak{p}}$ denote localization of $R$ and $M$ at $\mathfrak{p}$ respectively.
To prove this proposition we first set up some lemmasW.
lemma isLocalization_at_prime_prime_depth_le_depth [IsLocalRing Rₚ] [Module.Finite R M]
[ntr : Nontrivial Mₚ] : p.depth M IsLocalRing.depth M
This uses the regular-sequence characterization of depth. As mentioned in  2.4, the image of an $M$-regular sequence contained in $\mathfrak{p}$ under the map $R\to R_{\mathfrak{p}}$ is $M_{\mathfrak{p}}$-regular, yielding the above inequality.
Since we know $\operatorname{depth}(M_{\mathfrak{p}})\leq\dim(M_{\mathfrak{p}})$ from  3.3, to prove $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$, it remains to show $\operatorname{depth}_{\mathfrak{p}}(M)=\dim(M_{\mathfrak{p}})$W.
lemma isLocalize_at_prime_dim_eq_prime_depth_of_isCohenMacaulay
[Module.Finite R M] [M.IsCohenMacaulay] [ntr : Nontrivial Mₚ] :
Module.supportDim R M = p.depth M
From  3.3 we know that $\operatorname{depth}_{\mathfrak{p}}(M)\leq\operatorname{depth}(M_{\mathfrak{p}})$ is finite, so we proceed by induction on $\operatorname{depth}_{\mathfrak{p}}(M)$ following the proof in [4, Theorem 30(iii)].
Thus we finish proving that the localization at a prime of a Cohen–Macaulay module is Cohen–Macaulay, together with the equality $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$.
The formalization of a Cohen–Macaulay local ring is done more directly rather than by extending the module notionW.
class IsCohenMacaulayLocalRing : Prop extends IsLocalRing R where
depth_eq_dim : ringKrullDim R = IsLocalRing.depth (ModuleCat.of R R)
This is equivalent to (ModuleCat.of R R).IsCohenMacaulay when $R$ is a local ringW.
From the result that localization of a Cohen–Macaulay module at a prime is Cohen–Macaulay, we obtain the corresponding version for Cohen–Macaulay local ringsW.
lemma isCohenMacaulayLocalRing_localization_atPrime [IsCohenMacaulayLocalRing R] [IsNoetherianRing R]
(p : Ideal R) [p.IsPrime] (R : Type*) [CommRing Rₚ] [Algebra R Rₚ] [IsLocalization.AtPrime R p] :
IsCohenMacaulayLocalRing R
In a straightforward way, we define a Cohen–Macaulay ring as a ring whose localization at every prime is a Cohen–Macaulay local ringW.
class IsCohenMacaulayRing : Prop where
CM_localize : p : Ideal R, (_ : p.IsPrime), IsCohenMacaulayLocalRing (Localization.AtPrime p)
Using the above localization result for Cohen–Macaulay local rings, we formalized that the Cohen–Macaulay property need only be verified at maximal idealsW, and that a Noetherian ring is Cohen–Macaulay if and only if its localizations at maximal ideals are Cohen–Macaulay local ringsW W.
lemma isCohenMacaulayRing_iff [IsNoetherianRing R] : IsCohenMacaulayRing R
m : Ideal R, (_ : m.IsMaximal), IsCohenMacaulayLocalRing (Localization.AtPrime m)
lemma IsCohenMacaulayRing.of_isCohenMacaulayLocalRing [IsCohenMacaulayLocalRing R] [IsNoetherianRing R] : IsCohenMacaulayRing R
lemma IsCohenMacaulayLocalRing.of_isLocalRing_of_isCohenMacaulayRing [IsLocalRing R] [IsNoetherianRing R] [IsCohenMacaulayRing R] : IsCohenMacaulayLocalRing R
We also establish the corresponding results about quotienting a ring by a regular element or a regular sequenceW W.
lemma quotient_span_regular_isCohenMacaulay_iff_isCohenMacaulay [IsLocalRing R] [IsNoetherianRing R]
(x : R) (reg : IsSMulRegular R x) (mem : x maximalIdeal R) :
IsCohenMacaulayLocalRing R IsCohenMacaulayLocalRing (R Ideal.span {x})
lemma quotient_regular_sequence_isCohenMacaulay_iff_isCohenMacaulay [IsLocalRing R] [IsNoetherianRing R]
(rs : List R) (reg : IsWeaklyRegular R rs) (mem : r rs, r maximalIdeal R) :
IsCohenMacaulayLocalRing R IsCohenMacaulayLocalRing (R Ideal.ofList rs)
4.1. Unmixed Theorem for Cohen–Macaulay Rings
We first prove the following Theorem.
Theorem 4.1.
Let $R$ be a Cohen–Macaulay local ring and let $a_{1},\dots,a_{r}$ be a sequence in the unique maximal ideal $\mathfrak{m}$. The following are equivalent.
  1. (1)
    $a_{1},\cdots,a_{r}$ is $R$-regular
  2. (2)
    $0pt(a_{1},\cdots,a_{i})=i$, for all $1\leq i\leq r$
  3. (3)
    $0pt(a_{1},\cdots,a_{r})=r$
  4. (4)
    $a_{1},\cdots,a_{r}$ can be extended to a system of parameters.
[4, Theorem 31(iii)]
We do not state the theorem as a single List.TFAE, instead we provide step-by-step lemmas, since some parts do not essentially use the Cohen–Macaulay or even the local hypothesisW W W. In the following lemmas we assume $R$ is Noetherian.
lemma Ideal.ofList_height_eq_length_of_isWeaklyRegular (rs : List R) (reg : IsWeaklyRegular R rs)
(h : Ideal.ofList rs Top.top) : (Ideal.ofList rs).height = rs.length
lemma maximalIdeal_mem_ofList_append_minimalPrimes_of_ofList_height_eq_length [IsLocalRing R]
(rs : List R) (mem : r rs, r maximalIdeal R) (ht : (Ideal.ofList rs).height = rs.length) :
rs : List R, maximalIdeal R (Ideal.ofList (rs ++ rs’)).minimalPrimes
rs.length + rs’.length = ringKrullDim R
lemma isRegular_of_maximalIdeal_mem_ofList_minimalPrimes [IsCohenMacaulayLocalRing R]
(rs : List R) (mem : maximalIdeal R (Ideal.ofList rs).minimalPrimes)
(dim : rs.length = ringKrullDim R) : IsRegular R rs
The first two follow easily from Krull’s height theorem. The third follows from the fact that associated primes are minimal and that the quotient of a Cohen–Macaulay ring by a regular sequence is still Cohen–Macaulay, which we proved earlier in this section. The second and third together imply the following lemmaW.
lemma isRegular_of_ofList_height_eq_length_of_isCohenMacaulayLocalRing [IsCohenMacaulayLocalRing R]
(rs : List R) (mem : r rs, r maximalIdeal R) (ht : (Ideal.ofList rs).height = rs.length) :
IsRegular R rs
A by-product is the following lemmaW, i.e. [4, Theorem 31(i), Part1]:
lemma Ideal.depth_eq_height [IsCohenMacaulayLocalRing R] (I : Ideal R) (netop : I Top.top) :
I.depth (ModuleCat.of R R) = I.height
We then establish the following lemmaW to prove [4, Theorem 31(i), Part2]
lemma Ideal.primeHeight_add_ringKrullDim_quotient_eq_ringKrullDim [IsCohenMacaulayLocalRing R]
(p : Ideal R) [p.IsPrime] : p.primeHeight + ringKrullDim (R p) = ringKrullDim R
The formalization simply follows the proof in [4, Theorem 31(i), Part 2].
By simple manipulations of iSup and iInf we obtain the theoremW.
lemma Ideal.height_add_ringKrullDim_quotient_eq_ringKrullDim [IsCohenMacaulayLocalRing R]
(I : Ideal R) (netop : I Top.top) : I.height + ringKrullDim (R I) = ringKrullDim R
From this result and the fact that localization of a Cohen–Macaulay ring is still Cohen–Macaulay, for prime ideals $\mathfrak{q}\subseteq\mathfrak{p}$ we have $\dim(R_{\mathfrak{p}})=0pt(\mathfrak{q}R_{\mathfrak{p}})+\dim(R_{\mathfrak{p}}/\mathfrak{q}R_{\mathfrak{p}})$, i.e. $0pt(\mathfrak{p})-0pt(\mathfrak{q})=0pt(\mathfrak{p}/\mathfrak{q})$. This implies that a Cohen–Macaulay local ring is catenary. However, since notions of catenary order and catenary rings are not yet in mathlib, we do not develop this further here.
To prove the unmixed theorem, we first define the notion of an unmixed ideal as followsW.
class Ideal.IsUnmixed (I : Ideal R) : Prop where
height_eq : {p : Ideal R}, p associatedPrimes R (R I) p.height = I.height
With the results above, the formalization of the unmixed theorem follows easily by following the proof in [4, Theorem 32]W.
theorem isCohenMacaulayRing_iff_unmixed : IsCohenMacaulayRing R
(l : List R), (Ideal.ofList l).height = l.length (Ideal.ofList l).IsUnmixed
4.2. Universal Catenary
Using the results proven in  4.1, to prove that Cohen–Macaulay rings are universally catenary it remains only to show that a polynomial ring over a Cohen–Macaulay ring is also Cohen–MacaulayW.
theorem Polynomial.isCM_of_isCM [IsNoetherianRing R] [IsCohenMacaulayRing R] :
IsCohenMacaulayRing R[X]
Assume $R$ is Cohen–Macaulay. For any prime ideal $\mathfrak{p}$ of $R[X]$ set $\mathfrak{q}=\mathfrak{p}\cap R$. Note that $R[X]\mathfrak{p}$ is the localization of $R\mathfrak{q}[X]$ at a prime ideal, so the problem reduces to the following case: if $R$ is a Cohen–Macaulay local ring and $\mathfrak{p}$ is a prime ideal of $R[X]$ with $\mathfrak{p}\cap R$ equal to the maximal ideal of $R$, then $R[X]_{\mathfrak{p}}$ is a Cohen–Macaulay local ringW.
lemma Polynomial.localization_at_comap_maximal_isCM_isCM [IsNoetherianRing R]
[IsCohenMacaulayLocalRing R] (p : Ideal R[X]) [p.IsPrime] (max : p.comap C = maximalIdeal R) :
IsCohenMacaulayLocalRing (Localization.AtPrime p)
Using the flatness and localization results discussed in  2.4, the proof follows the argument in [4, Theorem 33].
Although the notions of catenary order and catenary rings are not yet available in mathlib, once they are added it will be straightforward to combine the above results to conclude that Cohen–Macaulay rings are universally catenary.
5. Regular Local Ring
Introduction:
In this section we first introduce our formalization of the definition of regular local rings and prove it is a domain. Then we prove the two directions of the Auslander–Buchsbaum–Serre criterion in the next two subsections. We first prove the global dimension of a regular local ring is equal to its Krull dimension via proving that a maximal Cohen–Macaulay module over a regular local ring is free. Then we prove the converse: if the unique maximal ideal of a Noetherian local ring has finite projective dimension then it is generated by a regular sequence; this finishes the converse and can be viewed as a weakened form of the Ferrand–Vasconcelos theorem. Finally, we establish some corollaries of the Auslander–Buchsbaum–Serre criterion in the last subsection, including that regularity of a ring can be checked at maximal ideals only, and Hilbert’s Syzygy theorem.
We give the definition of a regular local ring simply as a Noetherian local ring whose maximal ideal has span rank equal to its Krull dimensionW.
class IsRegularLocalRing : Prop extends IsLocalRing R, IsNoetherianRing R where
reg : (maximalIdeal R).spanFinrank = ringKrullDim R
Using the fact that the span rank of the maximal ideal equals the dimension of the cotangent space over the residue fieldW, we establish the equivalence of this definition with the cotangent-space formulationW.
lemma IsLocalRing.spanFinrank_maximalIdeal_eq_finrank_cotangentSpace [IsLocalRing R] [IsNoetherianRing R] :
(maximalIdeal R).spanFinrank = Module.finrank (ResidueField R) (CotangentSpace R)
lemma IsRegularLocalRing.iff_finrank_cotangentSpace [IsLocalRing R] [IsNoetherianRing R] :
IsRegularLocalRing R Module.finrank (ResidueField R) (CotangentSpace R) = ringKrullDim R
5.1. Regular Local Ring is Domain
To prove that a regular local ring is a domain and that minimal generators of the maximal ideal form a regular sequence, we establish the following lemmaW.
Lemma 5.1.
Let $(R,\mathfrak{m},k)$ be a regular local ring of dimension $d$, and let $a_{1},\cdots,a_{k}\in\mathfrak{m}$. Then the followings are equivalent.
  1. (1)
    $a_{1},\cdots,a_{k}$ can be extended to a regular system of parameters of $R$.
  2. (2)
    The images $\overline{a_{1}},\cdots,\overline{a_{k}}\in\mathfrak{m}/\mathfrak{m}^{2}$ are $k$-linearly independent.
  3. (3)
    $R/(a_{1},\cdots,a_{k})$ is a regular local ring of dimension dkd-k.
lemma quotient_isRegularLocalRing_tfae [IsRegularLocalRing R] (S : Finset R) (sub : (S : Set R) maximalIdeal R) :
[∃ (T : Finset R), S T T.card = ringKrullDim R Ideal.span T = maximalIdeal R,
LinearIndependent (ResidueField R) (((maximalIdeal R).toCotangent).comp (Set.inclusion sub)),
IsRegularLocalRing (R Ideal.span (S : Set R))
(ringKrullDim (R Ideal.span (S : Set R)) + S.card = ringKrullDim R)].TFAE
For $(1)\rightarrow(2)$: if $a_{1},\cdots,a_{k}$ extend to $a_{1},\cdots,a_{d}$ generating $\mathfrak{m}$, their images in the cotangent space are linearly independent, so $(2)$ holds. For $(2)\rightarrow(3)$: let $\tilde{\mathfrak{m}}$ the maximal ideal of $R/(a_{1},\cdots,a_{k})$. Then $\tilde{\mathfrak{m}}/\tilde{\mathfrak{m}}^{2}$ is isomorphic to $\mathfrak{m}/\mathfrak{m}^{2}/(\overline{a_{1}},\cdots,\overline{a_{k}})$, which has dimension dkd-k. Since the dimension of $R/(a_{1},\dots,a_{k})$ drops by at most $k$ from that of $R$, equality follows. For $(3)\rightarrow(1)$: pick generators $\tilde{a_{k+1}},\cdots,\tilde{a_{d}}$ of $\tilde{\mathfrak{m}}$ and lift them to $R$; together with $a_{1},\cdots,a_{k}$, these elements generate $\mathfrak{m}$, hence form a regular system of parameters.
Then we prove that a regular local ring $(R,\mathfrak{m},k)$ is a domain by induction on the dimensionW.
theorem isDomain_of_isRegularLocalRing [IsRegularLocalRing R] : IsDomain R
For dimension 0, $R$ is a field. Assume the statement proven for dimension $n$. For $n+1$, using prime avoidance applied to $\mathfrak{m}^{2}$ and all minimal primes, we obtain an element $x\in\mathfrak{m}$ not in $\mathfrak{m}^{2}$ not contained in any minimal prime. By the previous lemma, $R/(x)$ is a regular local ring of dimension $n$, hence by the induction hypothesis it is a domain, so $(x)$ is prime. As $x$ is not in any minimal prime, any minimal prime $\mathfrak{p}$ must satisfy $\mathfrak{p}\subsetneq(x)$, hence $\mathfrak{p}=x\cdot\mathfrak{p}$. Since $x\in\mathfrak{m}$ and $\mathfrak{p}$ is finitely generated, Nakayama’s lemma implies $\mathfrak{p}=0$. Thus $R$ is a domain.
For $R$ regular local ring of dimension $d$ and a sequence $rs$ of length $d$ generating its maximal ideal, any quotient of $R$ by sub-sequence of $rs$ is still regular and hence a domain, therefore $rs$ is a regular sequenceW.
theorem isRegular_of_span_eq_maximalIdeal [IsRegularLocalRing R] (rs : List R)
(span : Ideal.ofList rs = maximalIdeal R) (len : rs.length = ringKrullDim R) : IsRegular R rs
As a corollary, a regular local ring is Cohen–MacaulayW.
lemma isCohenMacaulayLocalRing_of_isRegularLocalRing [IsRegularLocalRing R] : IsCohenMacaulayLocalRing R
5.2. Global Dimension of Regular Local Ring
In this section, we prove that the global dimension of a regular local ring equals its Krull dimensionW.
theorem IsRegularLocalRing.globalDimension_eq_ringKrullDim [Small.{v} R] [IsRegularLocalRing R] :
globalDimension.{v} R = ringKrullDim R
Currently the definition of globalDimension depends on the universe we are in, further discussion appears in  6.1.
There is a well-known proof of this theorem using the following two facts for a Noetherian local ring $(R,\mathfrak{m},k)$.
  • $\operatorname{gldim}(R)=\operatorname{projdim}(k)$ [4, Theorem 41]
  • For any finitely generated $R$-module $M$ and $M$-regular element $x$, we have
    $\operatorname{projdim}(M/xM)=\operatorname{projdim}(M)+1$[4, Lemma 18.6]
Considering $k$ as the quotient of $R$ by a regular sequence finishes the proof. However, the $\operatorname{Tor}$ functor and related infrastructure are still incomplete in mathlib, so we cannot prove the first fact with the current setup. Instead we prove that maximal Cohen–Macaulay modules over a regular local ring are free, and use that to proceed.
5.2.1. Maximal Cohen–Macaulay Module
We first state a lemma used both here and in the next subsectionW.
Lemma 5.2.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring and $M$ a finitely generated $R$-module. Let $x\in\mathfrak{m}$ be $M$-regular. If $M/xM$ is free over $R/(x)$, then $M$ is free over $R$.
[6, 00NS]
lemma free_of_quotSMulTop_free [IsLocalRing R] [IsNoetherianRing R] (M : Type*) [AddCommGroup M]
[Module R M] [Module.Finite R M] {x : R} (mem : x maximalIdeal R) (reg : IsSMulRegular M x)
(free : Module.Free (R Ideal.span {x}) (QuotSMulTop x M)) : Module.Free R M
The formalization follows the argument in [6, 00NS].
We define a maximal Cohen–Macaulay module over $R$ simply as a module whose depth equals $\dim(R)$W.
class ModuleCat.IsMaximalCohenMacaulay [IsLocalRing R] [Small.{v} R] (M : ModuleCat.{v} R) : Prop where
depth_eq_dim : IsLocalRing.depth M = ringKrullDim R
The main result of this subsection is that a maximal Cohen–Macaulay module over a regular local ring is freeW. (See also [6, 00NT])
theorem free_of_isMaximalCohenMacaulay_of_isRegularLocalRing [IsRegularLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] [M.IsMaximalCohenMacaulay] : Module.Free R M
Proof sketch. We proceed by induction on $\dim(R)$. If $\dim(R)=0$, then $R$ is a field and $M$ is free. Assume the statement for $n$, and let $\dim(R)=n+1$. Choose $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$, by  5.1 the quotient $R/(x)$ is regular local ring of dimension $n$. To show $x$ is $M$-regular it suffices to prove that for every $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $x\notin\mathfrak{p}$. By the result in  3.3, $\dim(R/\mathfrak{p})\geq\operatorname{depth}(M)=\dim(R)=n+1$. If $x\in\mathfrak{p}$ then $n=\dim(R/(x))\geq\dim(R/\mathfrak{p})$, causing a contradiction. Hence $x$ is $M$-regular. Using results in  3.1 we get $\operatorname{depth}(M/xM)=\operatorname{depth}(M)-1=\dim(R)-1=n=\dim(R/(x))$, so $M/xM$ is maximal Cohen–Macaulay over $R/(x)$. By the induction hypothesis, $M/xM$ is free over $R/(x)$, then the result follows from  5.2.
5.2.2. The Main Proof
The proof idea is standard. For regular local ring $(R,\mathfrak{m},k)$ of dimension $d$ and any $R$-module $M$, repeated application of  3.2 shows that the $d$-th syzygy of $M$ in a free resolution is either 0 or maximal Cohen–Macaulay, in either case it is projective. (See [1, Theorem 2.2.7]) However, formalizing a distant syzygy and the depth calculations is cumbersome, so we use a trick to simplify the formalization.
Results in  3.2 imply that for a finitely generated $R$-module $M$ of finite projective dimension, the projective dimension does not exceed $\operatorname{depth}(R)=\dim(R)$. Since $\operatorname{depth}(k)=0$, if $k$ has finite projective dimension then its projective dimension equals $\operatorname{depth}(R)=\dim(R)$. Thus it suffices to show every finitely generated module over a regular local ring has finite projective dimension W.
lemma projectiveDimension_ne_top_of_isRegularLocalRing [IsRegularLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] : projectiveDimension M Top.top
We set up an auxiliary lemma that is easier to prove by induction on an integer $i$, using only single-step dimension shiftingW.
lemma finite_projectiveDimension_of_isRegularLocalRing_aux [IsRegularLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] (i : ℕ) :
IsLocalRing.depth M + i ringKrullDim R n, HasProjectiveDimensionLE M n
For $i=0$, $\operatorname{depth}(M)\geq\dim(R)$ implies equality, so $M$ is maximal Cohen–Macaulay and hence free by the previous lemma; thus $M$ has finite projective dimension. Assume the claim for $i$. For $i+1$, if $M=0$ the claim is trivial. Otherwise take a short exact sequence $0\to N\to F\to M\to 0$ with $F$ free and finitely generated. By  3.2, $\operatorname{depth}(N)+i\geq min{\operatorname{depth}(F),\operatorname{depth}(M)+1}+i=min{\operatorname{depth}(F)+i,\operatorname{depth}(M)+i+1}$. Since $\operatorname{depth}(F)=\operatorname{depth}(R)=\dim(R)$, the right-hand side is at least $\dim(R)$, so $\operatorname{depth}(N)+i\geq\dim(R)$. By the induction hypothesis $N$ has finite projective dimension, hence so does $M$. This completes the induction.
Combining the above results yields the formal proof of the theorem stated at the beginning of this subsection.
5.3. A Weakened Version of Ferrand–Vasconcelos Theorem
We first establish a lemma about projective dimension: for a finitely generated module $M$ over a Noetherian local ring $R$, and for $x$ that is both $R$-regular and $M$-regular, we have $\operatorname{projdim}_{R}(M)=\operatorname{projdim}_{R/(x)}(M/xM)$W. (See [1, Lemma 1.3.5].)
lemma projectiveDimension_eq_quotient [Small.{v} R] [IsLocalRing R] [IsNoetherianRing R]
(M : ModuleCat.{v} R) [Module.Finite R M]
(x : R) (reg1 : IsSMulRegular R x) (reg2 : IsSMulRegular M x) (mem : x maximalIdeal R) :
projectiveDimension.{v} M = projectiveDimension.{v} (ModuleCat.of (R Ideal.span {x}) (QuotSMulTop x M))
Since projective and free coincide over a local ring, Lemma  5.2 implies that $M$ is projective over $R$ if and only if $M/xM$ is projective over $R/(x)$. Using dimension shifting, it suffices to reduce the general equality of projective dimensions to the equivalence of projectivity on both sides. Concretely, take a short exact sequence $0\rightarrow N\xrightarrow{f}R^{n}\xrightarrow{g}M\rightarrow 0$. Tensoring with $R/(x)$ yields an exact sequence $N/xN\rightarrow\rightarrow(R/(x))^{n}\rightarrow M/xM\rightarrow 0$ with the obvious transition map. Any element $y$ in the kernel of the map $N\rightarrow R^{n}\rightarrow(R/(x))^{n}$ lifts to an element of $N$ of the form $y=x\cdot z$. Since $x\cdot g(z)=g(x\cdot z)=0$, we have $z\in N$. Thus the kernel of $N\rightarrow N/xN\rightarrow(R/(x))^{n}$ equals $xN$, so the induced map $N/xN\rightarrow(R/(x))^{n}$ is injective. Because $\operatorname{projdim}_{R}(N)=\operatorname{projdim}_{R}(M)-1$ and $\operatorname{projdim}_{R/(x)}\ N/xN=\operatorname{projdim}_{R/(x)}\ M/xM-1$, and since $x$ remains $R$-regular and $N$-regular (as $N$ is submodule of a free module), the equality of projective dimensions follows. (The exactness is essentially a computation of $\operatorname{Tor}_{R}^{1}(R/(x),M)$)
Returning to the main thread, the Ferrand–Vasconcelos theorem states: for an ideal $I$ of a Noetherian local ring $R$, if $I$ has finite projective dimension and $I/I^{2}$ is free over $R/I$, then $I$ is generated by a regular sequence. (See [1, Theorem 2.2.8]) If the residue field $k$ of a Noetherian local ring $(R,\mathfrak{m},k)$, has finite projective dimension, then so does $\mathfrak{m}$; applying the theorem to $\mathfrak{m}$ shows that $\mathfrak{m}$ is generated by a regular sequence. Since any regular sequence contained in $\mathfrak{m}$ has length at most $\operatorname{depth}(R)\leq\dim(R)$, this implies that $R$ is regular.
However, proving the full Ferrand–Vasconcelos theorem requires showing that any ideal $I$ with a finite free resolution contains a $R$-regular element (See [1, Corollary 1.4.7]), which in turn requires a theory of rank for free resolutions. As the theory of rank in mathlib is not yet fully developed, the complete proof is currently out of reach. The key observation is that the induction in the Ferrand–Vasconcelos proof still works if we only consider the maximal ideal. Thus, we restrict our attention to the unique maximal ideal and prove the result for that case; this already yields regularity provided the residue field has finite projective dimensionW.
theorem generate_by_regular [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R]
(h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
rs : List R, IsRegular R rs Ideal.ofList rs = maximalIdeal R
The existence of a regular element in the maximal ideal, assuming the maximal ideal has finite projective dimension and $R$ is not a field, follows from the Auslander–Buchsbaum equality (see  3.2)W.
lemma exist_isSMulRegular_of_exist_hasProjectiveDimensionLE_aux [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (nebot : maximalIdeal R ⊥)
(h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
x maximalIdeal R, IsSMulRegular R x
Proof sketch. Suppose there are no regular elements in $\mathfrak{m}$. Then $\operatorname{depth}(R)=0$. If $k$ has finite projective dimension, the Auslander–Buchsbaum formula (see  3.2) gives $\operatorname{projdim}(k)+\operatorname{depth}(k)=\operatorname{depth}(R)=0$. Since $\operatorname{depth}(k)=0$, we get $\operatorname{projdim}(k)=0$, so $k$ is projective hence free; thus $\mathfrak{m}=0$,contradicting the assumption that $R$ is not a field.
Using prime avoidance applied to $\mathfrak{m}^{2}$ and the associated primes of $R$, we can further find a $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$W.
lemma exist_isSMulRegular_of_exist_hasProjectiveDimensionLE [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (nebot : maximalIdeal R ⊥)
(h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
x maximalIdeal R, x maximalIdeal R ^ 2 IsSMulRegular R x
We proceed by induction on the minimal number of generators (span rank) of $\mathfrak{m}$. If this number is 0, $\mathfrak{m}=0$ is generated by the empty sequence. Assume the statement for $n$, and consider the case $n+1$. Choose a $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ as above. Denote by $\tilde{\mathfrak{m}}$ the maximal ideal of $R/(x)$, its span rank is exactly $n$. To apply the induction hypothesis to $R/(x)$, it remains to show $\tilde{\mathfrak{m}}$ has finite projective dimension over $R/(x)$. Let $\overline{J}$ be a direct sum complement of $\mathrm{span}\{\overline{x}\}$ in $\mathfrak{m}/\mathfrak{m}^{2}$, and let $J$ be the preimage of $\overline{J}$ in $\mathfrak{m}$. Then $J+(x)=\mathfrak{m}$. For any $y\in J\cap(x)$, write $y=xz$. Because $\overline{y}\in\mathrm{span}\{\overline{x}\}\cap\overline{J}=0$ in $\mathfrak{m}/\mathfrak{m}^{2}$, we get $z\in\mathfrak{m}$ and hence $y\in x\mathfrak{m}$. Thus $J\cap(x)\subseteq x\mathfrak{m}$. From this one obtains maps whose composition is the identity:
$\tilde{\mathfrak{m}}\cong(J+(x))/(x)\cong J/(J\cap(x))\rightarrow\mathfrak{m}/x\mathfrak{m}\rightarrow\tilde{\mathfrak{m}}$
Therefore $\tilde{\mathfrak{m}}$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$ as a $R/(x)$ module. By the first lemma of this section, $\operatorname{projdim}_{R/(x)}\ \mathfrak{m}/x\mathfrak{m}=\operatorname{projdim}_{R}\mathfrak{m}<\infty$, so $\tilde{\mathfrak{m}}$ has finite projective dimension. Applying the induction hypothesis, $\tilde{\mathfrak{m}}$ is generated by a $R/(x)$-regular sequence $rs$. Prepending $x$ yields $x::rs$, a $R$-regular sequence generating $\mathfrak{m}$. This completes the induction.
We conclude with the following two theoremsWW.
theorem IsRegularLocalRing.of_maximalIdeal_hasProjectiveDimensionLE [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R]
(h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
IsRegularLocalRing R
theorem IsRegularLocalRing.of_globalDimension_lt_top [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (h : globalDimension.{v} R < Top.top) : IsRegularLocalRing R
5.4. The Final Conclusions and Related Results
5.4.1. Auslander–Buchsbaum–Serre Criterion
With the results in  5.2 and  5.3, we can finally state the Auslander–Buchsbaum–Serre criterion as followsW.
theorem Auslander_Buchsbaum_Serre [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R] :
IsRegularLocalRing R globalDimension.{v} R < Top.top
As noted in  2.5, global dimension of a ring $R$ equals the supremum of the global dimensions of the localizations $R_{\mathfrak{p}}$ as $\mathfrak{p}$ ranges over the prime ideals. Hence if $R$ has finite global dimension then each $R_{\mathfrak{p}}$ does as well. A useful corollary of the Auslander–Buchsbaum–Serre criterion is that the localization of a regular local ring at any prime is again regularW.
lemma IsRegularLocalRing.of_isLocalization [IsRegularLocalRing R] (p : Ideal R) [p.IsPrime]
(S : Type*) [CommRing S] [Algebra R S] [IsLocalization.AtPrime S p] : IsRegularLocalRing S
5.4.2. Regular Ring
We also developed basic theory of regular rings. The definition of a regular ring is simply that its localization at every prime is a regular local ringW.
class IsRegularRing : Prop where
localization_isRegular : p : Ideal R, (_ : p.IsPrime),
IsRegularLocalRing (Localization.AtPrime p)
Using  2.5 and the fact that a regular local ring has global dimension equal to its Krull dimension, we obtain the analogous statement for regular ringsW.
theorem IsRegularRing.globalDimension_eq_ringKrullDim [Small.{v} R] [IsRegularRing R] [IsNoetherianRing R] :
globalDimension.{v} R = ringKrullDim R
Since localization of a regular local ring at a prime is regular, regularity of a ring can be verified on maximal ideals onlyW.
lemma isRegularRing_of_localization_maximal_isRegularLocalRing
(h : m : Ideal R, (_ : m.IsMaximal), IsRegularLocalRing (Localization.AtPrime m)) :
IsRegularRing R
Finally, we prove that a regular local ring is a regular ringW.
lemma isRegularRing_of_isRegularLocalRing [IsRegularLocalRing R] : IsRegularRing R
5.4.3. Hilbert’s Syzygy Theorem
To prove Hilbert’s Syzygy Theorem (see [4, Corollary 18.1]), it remains to show that a polynomial ring over a regular ring is regularW.
theorem Polynomial.isRegularRing_of_isRegularRing [IsRegularRing R] :
IsRegularRing R[X]
Because both regularity and the Cohen–Macaulay property are local conditions, the proof reduces similarly to the following local case: for a regular local ring $(R,\mathfrak{m},k)$ and a prime ideal $\mathfrak{p}$ of $R[X]$ with $\mathfrak{p}\cap R=\mathfrak{m}$, show that $R[X]_{\mathfrak{p}}$ is a regular local ringW.
lemma Polynomial.localization_at_comap_maximal_isRegularRing_isRegularRing
[IsRegularLocalRing R] (p : Ideal R[X]) [p.IsPrime] (max : p.comap C = maximalIdeal R) :
IsRegularLocalRing (Localization.AtPrime p)
One shows that $\mathfrak{p}$ can be generated by $0pt(\mathfrak{p})=\dim(R[X]_{\mathfrak{p}})$ elements, these generators then generate the maximal ideal of $R[X]_{\mathfrak{p}}$. If $\mathfrak{p}=\mathfrak{m}[X]$, it is generated by the minimal generators of $\mathfrak{m}$. Since $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})=\dim(R)$, the localization $R[X]_{\mathfrak{p}}$ is regular. If $\mathfrak{m}[X]\subsetneq p$, the image of $\mathfrak{p}$ in $k[X]$ a single polynomial $\overline{f}$. Lift $\overline{f}$ to $f\in R[X]$, then $\mathfrak{p}$ is generated by $f$ together with the minimal generators of $\mathfrak{m}$. As $0pt(\mathfrak{p})=0pt(\mathfrak{m})+1=\dim(R)+1$, the localization $R[X]_{\mathfrak{p}}$ is regular in this case as well. This completes the proof.
Thus we formalize Hilbert’s Syzygy theorem as followsW.
theorem Hilberts_Syzygy (k : Type u) [Field k] [Small.{v} k] (n : ℕ) :
globalDimension.{v} (MvPolynomial (Fin n) k) = n
6. Discussion
6.1. Remarks About Implementation
In formalization of category theory in Lean4, one often needs to take universes into careful consideration. For example, in this project we must specify the universe in which the $\operatorname{Ext}$ functor lives. In our application we mainly use the instance for having $\operatorname{Ext}$ for categories with enough projectives W. For the category of $R$-modules in universe $v$ with $R$ in universe $u$, if we assume $R$ is $v$-small (denoted [Small.{v} R]), then we can obtain $\operatorname{Ext}$ in any universe greater than or equal to $v$. In our implementation, we only explicitly specify the universe level of $\operatorname{Ext}$ when it actually appears in the statement of a theorem; otherwise we hide it or fix a universe level. For example, for projectiveDimension and globalDimension, we use the standard universe following previous definitionsW. For the definition of depth, with $R$ in universe u and assuming Small.{v} R, in ModuleCat.{v} R we consider $\operatorname{Ext}$ in universe v, which is the smallest universe in which we obtain $\operatorname{Ext}$ using CategoryTheory.hasExt_of_enoughProjectives W. None of this matters much in practice, since there are already definitions for changing universes W, which simply pass through shifted hom, moreover, if the category is $R$-linear, then this is an $R$-linear equivalence and the additional structure is preserved.
A remaining nuisance is the universe of the modules on which we take $\operatorname{Ext}$. Since the commutation of $\operatorname{Ext}$ and Ulift is not fully formalized yet, some definitions still partially depend on a choice of universe. However, we know that $\operatorname{Ext}$ and Ulift do commute, so this will be resolved once that fact is formalized.
6.2. Related Works
As far as we are aware, in other theorem provers the concepts of derived categories and derived functors are often not formalized, let alone theories built on them such as $\operatorname{depth}$ and homological dimension. In Lean4, this work is also a first step toward applications of homological methods in commutative algebra.
For other related formalizations of commutative and homological algebra in Lean4, as mentioned in  2.3, there is a project that formalized that the Krull dimension of a polynomial ring over a Noetherian ring increases by exactly one (seeW). That proof proceeds by showing (a version of) [4, Theorem 19] : if $R\rightarrow S$ satisfies going down property, then the height of a prime ideal $P$ of $S$ equals the height of its preimage $\mathfrak{p}$ in $R$ plus the height of the image of $P$ in $S/\mathfrak{p}S$ (seeW). All the results about derived categories and $\operatorname{Ext}$ in mathlib that we use in this project originate from the Liquid Tensor Experiment development [2] and were ported to Lean4 [5].
6.3. Future Works
Building on the development of Cohen–Macaulay rings and regular local rings, there are several natural directions for future work.
Cohen–Macaulay Modules over Arbitrary Ring
For a non‑local ring $R$, one can define Cohen–Macaulay modules as follows: an $R$-module $M$ is Cohen–Macaulay if for every prime ideal $\mathfrak{p}$ of $R$, the localized module $M_{\mathfrak{p}}$ is Cohen–Macaulay over $R_{\mathfrak{p}}$.
Miracle Flatness
For a local homomorphism $R\to S$ with $R$ a regular local ring and $S$ a Cohen–Macaulay local ring, if $\dim S=\dim R+\dim(S/\mathfrak{m}_{R}S)$, then $R\to S$ is flat. This statement is known as miracle flatness (see [6, 00R4]). After the developments in this project, most of the required techniques are already formalized or are only a small step away. The remaining obstacle is formalizing a variant of the local criterion for flatness:
Lemma 6.1.
Let $R\rightarrow S$ be a local homomorphism of Noetherian local rings. Let $I\neq R$ be an ideal in R. Let M be a finite $S$-module. If $\operatorname{Tor}_{R}^{1}(M,R/I)=0$ and $M/IM$ is flat over $R/I$, then $M$ is flat over $R$.
[6, 00ML]
Criterion of Normality
We use the following standard notation for a ring $R$.
  • ($S_{k}$) : $\operatorname{depth}(R_{\mathfrak{p}})\geq min\{k,0pt(\mathfrak{p})\}$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$
  • ($R_{k}$) : if $\mathfrak{p}\in\operatorname{Spec}(R)$ and $0pt(\mathfrak{p})\leq k$, $R_{\mathfrak{p}}$ is regular.
Given our formalization of depth, we can state the criterion of normality as follows: for a Noetherian ring $R$, $R$ is normal iff it satisfies ($S_{2}$) and ($R_{1}$) (see [4, Theorem 39]). Formalizing this criterion would be an interesting project that will develop many results about normal domains along the way.