### v3

# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

Abstract.
We present a comprehensive formalization in the Lean4 theorem prover of the Auslander–Buchsbaum–Serre criterion, which characterizes regular local rings as those Noetherian local rings with finite global dimension. Rather than following the well-known proof that computes the projective dimension of the residue field via quotient by regular sequences and uses the Koszul complex to bound the cotangent space dimension by the global dimension, our approach is built systematically on the formalization of depth defined via the vanishing of Ext functors. We establish key homological results including Rees’ theorem, the Auslander–Buchsbaum formula, and Ischebeck’s theorem, and further develop the theories of Cohen–Macaulay modules and rings, including a complete formalization of the unmixedness theorem for Cohen–Macaulay rings. To prove the Auslander–Buchsbaum–Serre criterion, we show that maximal Cohen–Macaulay modules over regular local rings are free and establish a weakened form of the Ferrand–Vasconcelos theorem specific for the unique maximal ideal. As corollaries, we deduce that regularity can be checked at maximal ideals and formalize Hilbert’s Syzygy Theorem. This work demonstrates how homological algebra can be effectively employed in the formalization of commutative algebra, providing extensive infrastructure for future developments in the field.
1. Introduction
The Auslander–Buchsbaum–Serre criterion is a beautiful mathematical result stating that for a Noetherian local ring $R$, the ring $R$ is regular if and only if it has finite global dimension. This criterion was not proved all at once. At first, only the “only if” direction was known, as shown in [3]. At that time, it was still an open problem whether the localization of a regular local ring at a prime ideal remains regular. The homological characterization of regular local rings was completed only after Jean–Pierre Serre proved the converse direction later in [20], as formally recorded in [1]. Once this characterization is available, it becomes straightforward to prove that the localization of a regular local ring is again regular. This is one of the earliest examples where homological methods were used to prove a result that is not itself homological. It shows how homological invariants—such as projective dimension, depth, and homological dimension—can detect and control purely “geometric” or “structural” properties of rings. This opened the door for homological algebra to become a central tool in commutative algebra and algebraic geometry. A similar homological perspective later led to the Auslander–Buchsbaum formula, Serre’s intersection multiplicity theory [21], the development of Cohen–Macaulay and Gorenstein ring theory, and deformation theory, for example [9].
This project began with the goal of developing the theories of depth and Cohen–Macaulay rings. As the concepts of regular element and regular sequence were already formalized, this goal was within reach using the current library mathlib [15]. The effort was not only of mathematical interest, but also a test of the newly developed homological algebra in mathlib via its application to commutative algebra. As we shifted our focus toward combining it with homological algebra, we set Hilbert’s Syzygy theorem and the Auslander–Buchsbaum–Serre criterion as our main targets. However, at the time of this work, the $\operatorname{Tor}$ functor was still largely missing from mathlib, and the development of the Koszul complex was also incomplete, so the most well-known proof could not be used in the formalization. The turning point was finding an approach to the global dimension of a regular local ring by proving that a maximal Cohen–Macaulay module over a regular local ring is free; then the target became attainable. Later, inspired by the proof of the Ferrand–Vasconcelos theorem in [5, Theorem 2.2.8], we realized that induction on span rank (minimal number of generators) could still be carried out by focusing only on the unique maximal ideal. This allowed us to complete the full Auslander–Buchsbaum–Serre criterion. The value of this work lies not only in the final theorems: throughout the process, we formalized more basic constructions in commutative and homological algebra (for example, the definitions of projective dimension and $\operatorname{depth}$), tested these definitions in later developments, and improved their interfaces to make them easier to use. However, some of these additional interfaces are not included in this paper due to space limitations. Currently part of subsection  2.5 and all of subsections  2.1,  2.2,  2.4 are already in mathlib. We plan to upstream most of our work into mathlib in the future.
Outline of the Paper
In Section  2, we introduce the preliminary results needed for the formalization, including existing theories in mathlib and supplementary results we developed during the study of Cohen–Macaulay rings and regular local rings. We then present the development of depth and Cohen–Macaulay modules and rings in Sections  3 and  4, following [16, Section 6]. This was the earliest part to be formalized.
The key preliminary ingredient of the Auslander–Buchsbaum–Serre criterion is the Auslander–Buchsbaum formula (Theorem  3.8) in Subsection  3.2.
Theorem 1.1 (Auslander–Buchsbaum).
Let $(R,\mathfrak{m})$ be a Noetherian local ring and let $M$ be a non-trivial finitely generated $R$-module. If projective dimension of $M$ is finite, then
$\operatorname{projdim}(M)+\operatorname{depth}(M)=\operatorname{depth}(R)$
After developing the basic theory of regular local rings at the beginning of Section  5, and establishing that a regular local ring is a domain in Subsection  5.1, the proof of the criterion splits into two largely independent parts, corresponding to the two directions of the equivalence.
  • In subsection  5.2, we prove that a regular local ring has global dimension equal to its Krull dimension, which establishes one direction of the criterion.
Theorem 1.2.
For regular local ring $R$, $\operatorname{gldim}(R)=\dim(R)$.
Let $(R,\mathfrak{m},k)$ be a regular local ring of dimension $d$. By Theorem  1.1, the projective dimension of any finitely generated module is either infinite or at most $\operatorname{depth}(R)=d$. Since $\operatorname{depth}(k)=0$, if $k$ has finite projective dimension, then again by Theorem  1.1, we have $\operatorname{projdim}(k)=d$. This reduces the problem to showing that all finitely generated modules over a regular local ring have finite projective dimension.
By basic inequalities of $\operatorname{depth}$, we obtain the following:
Definition 1.3.
A module $M$ over local ring $R$ is a maximal Cohen–Macaulay module if $\operatorname{depth}(M)=\dim(R)$.
Proposition 1.4.
The $d$-th syzygy of a finitely generated module $M$ in a free resolution is either a maximal Cohen–Macaulay module or 0
The next result about maximal Cohen–Macaulay modules then ensures that this yields a finite projective (free) resolution of $M$.
Theorem 1.5 ([22, 00NT]).
For regular local ring $R$, any finitely generated maximal Cohen–Macaulay module over $R$ is free.
If an element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ is $M$-regular, then $M/xM$ is a maximal Cohen–Macaulay module over $R/(x)$ by the following proposition. This allows Theorem  1.5 to be proved by induction on $\dim(R)$.
Proposition 1.6.
For $R$-modules $M$ and $N$ and an $M$-regular element $x\in\mathrm{Ann}(N)$, we have
$\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$
  • For the converse direction–showing regularity from finite global dimension, we initially attempted to use the following theorem (see Subsection  5.3).
Theorem 1.7 (Ferrand–Vasconcelos, [5, Theorem 2.2.8]).
For an ideal $I$ of a Noetherian local ring $R$, if $I$ has finite projective dimension and $I/I^{2}$ is free over $R/I$, then $I$ is generated by a regular sequence.
However, since some techniques in its proof are not yet fully formalized, we focus only on the case where $I$ is the unique maximal ideal, which still suffices for our goal. Note that a regular sequence in $\mathfrak{m}$ has length at most $\operatorname{depth}(R)\leq\dim(R)$. If $\mathfrak{m}$ has finite projective dimension, the above theorem directly implies that $R$ is regular.
Returning to the case $I=\mathfrak{m}$, the proof roughly follows a modified version combining ideas from [22] and [5, Section 2.2]. By induction on the minimal number of generators of $\mathfrak{m}$, if $x$ is a regular element in $\mathfrak{m}\setminus\mathfrak{m}^{2}$, then $\mathfrak{m}/(x)$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$. The following proposition then implies that $\mathfrak{m}/(x)$ has finite projective dimension, so the induction hypothesis applied to $R/(x)$ completes the proof.
Proposition 1.8.
For $R$-module $M$ and an element $x$ being both $R$-regular and $M$-regular, then
$\operatorname{projdim}_{R/(x)}(M/xM)=\operatorname{projdim}_{R}(M)$
It is worth noting that many properties of Cohen–Macaulay rings are established independently, and not all results presented here are directly tied to the Auslander–Buchsbaum–Serre criterion. However, most of these unrelated results turn out to be useful in proving the unmixedness theorem for Cohen–Macaulay rings.
Theorem 1.9.
A Noetherian ring $R$ is Cohen–Macaulay if and only if for any ideal $I=(r_{1},\cdots,r_{k})$ with $0pt(I)=k$, II is unmixed.
In Subsection  5.4, we state our final results and their corollaries, including another meaningful result obtained along the way: Hilbert’s Syzygy Theorem, which only requires the regularity of the polynomial ring in addition to our formalization.
Theorem 1.10 (Hilbert’s Syzygy).
For a field $k$, $\operatorname{gldim}(k[x_{1},\cdots,x_{n}])=n$.
Finally, in section  6, we describe implementation issues–particularly the universe issues, related works, and directions for future works.
Throughout the paper, we use the symbol W for external links. Most statements and definitions are accompanied by such a link directly to the source code for the corresponding statement in mathlib or in a branch of our fork W. To keep the links stable, the links to the latter point to a fixed commit of the master branch; while other links point to the documentation of mathlib. In some code excerpts the dot notation is exhibited: for example, for I of type Ideal R, the call Ideal.depth I M can be shortened to I.depth M.
Acknowledgements
We thank Professor Liang Xiao for his preliminary review and for organizing the collaborative work that made this project possible. We are grateful to Dr. Jiedong Jiang for his helpful suggestions in revising the first version into the second. We also thank Dr. Huanhuan Yu for producing the blueprint of the formalization project. We are also deeply grateful to all the contributors and reviewers who have been involved in the development and update process of both Lean and mathlib, whose efforts have made our work possible. Special thanks go to Professor Joël Riou for developing the derived category and the $\operatorname{Ext}$ functor, which enabled the homological methods. This work is supported in part by National Key R&D Program of China grant 2024YFA1014000.
2. Preliminaries
In this section, we review some preliminary results that will be used in later proofs. All code shown in this section are due to the authors, except for those indicated in Subsections  2.1 and  2.5.
2.1. Associated Primes
At the start of this project, we only had the definition associatedPrimes W and basic facts about associated primes in mathlibW. Thanks to the work of Jinzhao Pan, we now have the finiteness of associated primes for finitely generated modules over Noetherian rings in mathlibW.
theorem associatedPrimes.finite (A : Type u) [CommRing A] (M : Type v) [AddCommGroup M] [Module A M]
[IsNoetherianRing A] [Module.Finite A M] : (associatedPrimes A M).Finite
Combined with the fact that the complement of the non-zero-divisors of a module equals the union of its associated primesW, this lemma allows us to apply the prime avoidance lemma when looking for a regular element in an ideal.
In addition to existing results in mathlib, we formalized results on associated primes of localized modules.W
Lemma 2.1.
For $S$ a multiplicative set of a Noetherian ring $R$ and $M$ a finitely generated $R$-module, one has $\mathrm{Ass}(S^{-1}M)=\mathrm{Ass}(M)\cap\operatorname{Spec}(S^{-1}R)$.
Denoting $R^{\prime}$ and $M^{\prime}$ as the localizations of $R$ and $M$ at $S$ respectively, the main statement is:
lemma preimage_comap_associatedPrimes_eq_associatedPrimes_of_isLocalizedModule [IsNoetherianRing R] :
(Ideal.comap (algebraMap R R’)) ⁻¹’ (associatedPrimes R M) = associatedPrimes R M
Since the induced map $\operatorname{Spec}(R^{\prime})\to\operatorname{Spec}(R)$ is injective, this gives a full characterization of $\mathrm{Ass}(S^{-1}M)$.
Furthermore, applying the above to the localization at $p$, a minimal prime over $\mathrm{Ann}(M)$, since associated primes of $M_{p}$ exist and any such associated prime must be the maximal ideal of $R_{p}$ (because $p$ is minimal), we obtain the following lemma.W
Lemma 2.2.
For $M$ a finitely generated module over Noetherian ring $R$, minimal primes over $\mathrm{Ann}(M)$ lie in $\mathrm{Ass}(M)$.
lemma minimalPrimes_annihilator_subset_associatedPrimes [IsNoetherianRing R] [Module.Finite R M] :
(Module.annihilator R M).minimalPrimes associatedPrimes R M
2.2. Krull Dimension of Module
Recall that the Krull dimension of a module is the Krull dimension of its support viewed as a preordered set. We implement it in Lean4 in the same way.W
def supportDim : WithBot ℕ∞ := krullDim (Module.support R M)
Some basic results are formalized hereW.
We prove a lemma that will be frequently used later.W
Lemma 2.3.
For a finitely generated $R$-module $M$ and an $M$-regular element $x$ contained in the Jacobson radical of $\mathrm{Ann}(M)$, then $\dim(M/xM)+1=\dim(M)$.
theorem supportDim_quotSMulTop_succ_eq_supportDim_mem_jacobson {x : R} (reg : IsSMulRegular M x)
(hx : x (annihilator R M).jacobson) : supportDim R (QuotSMulTop x M) + 1 = supportDim R M
Proof.
Note that any $M$-regular element is not contained in any minimal prime over $\mathrm{Ann}(M)$. The following two propositions suffices.
Proposition 2.4.
For a finitely generated $R$-module $M$, quotienting by an element in the Jacobson radical of $\mathrm{Ann}(M)$ decreases the dimension by at most one.
Proof.
The formalization of this proposition follows [10, Proposition 2.3.4] via [10, Lemma 2.3.3]W. ∎
Proposition 2.5.
For a finitely generated $R$-module $M$, quotienting by an element decreases the dimension by at least one.
By induction on the length of a regular sequence, we obtain the following corollary of Lemma  2.3.W
Corollary 2.6.
If $R$ is local, for a finitely generated $R$-module $M$ and an $M$-regular sequence $(r_{1},\cdots,r_{k})$, $\dim(M/(r_{1},\cdots,r_{k})M)+k=\dim(M)$.
theorem supportDim_add_length_eq_supportDim_of_isRegular (rs : List R) (reg : IsRegular M rs) :
supportDim R (M ofList rs (Top.top : Submodule R M)) + rs.length = supportDim R M
2.3. Dimension of Polynomial Rings
As we need to work with polynomials over Cohen–Macaulay and regular rings, some results about the heights of ideals in polynomial rings over Noetherian rings are required. Let $R$ be a Noetherian ring. We establish the following results, which follow easily from Krull’s height theoremW.
Proposition 2.7.
For a local ring $(R,\mathfrak{m})$, $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})$
Proposition 2.8.
For a ring $R$, let $\mathfrak{p}$ be a prime ideal of $R[X]$, then $0pt(\mathfrak{p})\leq 0pt(\mathfrak{p}\cap R)+1$
From these we obtain the standard result on the Krull dimension of polynomial rings over Noetherian rings.W
lemma ringKrullDim_of_isNoetherianRing [IsNoetherianRing R] : ringKrullDim R[X] = ringKrullDim R + 1
There is an ongoing effort to formalize similar statements more generally by studying heights under ring homomorphisms with the going-down property.WW This would establish more general versions of the above ad hoc lemmas. These lemmas will be deprecated once the generalizations become available in mathlib. See Subsection  6.2 for details.
2.4. Regular Sequence Under Flat Base Change
Since we need to handle preservation of regular sequences under localization, we formalized the more general results that weakly regular sequences are preserved under flat base changeW and regular sequences are preserved under faithfully flat base changeW. These results also benefit the proof that polynomials over Cohen–Macaulay rings are again Cohen–Macaulay in Subsection  4.2.
theorem IsWeaklyRegular.of_flat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S]
{M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N]
[IsScalarTower R S N] [Module.Flat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R}
(reg : IsWeaklyRegular M rs) : IsWeaklyRegular N (List.map (algebraMap R S) rs)
theorem IsRegular.of_faithfullyFlat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S]
{M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N]
[IsScalarTower R S N] [Module.FaithfullyFlat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R}
(reg : IsRegular M rs) : IsRegular N (List.map (algebraMap R S) rs)
Specializing to the case where the module is the ring itself yields a more user-friendly version.WW
Finally, we also added specialized lemmas for localized modules and localization, which show that a regular sequence contained in the prime being localized at remains regular after localization.WW
2.5. Projective Dimension
Since the $\operatorname{Ext}$ functor was introduced to mathlib relatively recently as presented in [19], at the time of this work, besides the basic constructions inherited from shifted hom in derived category such as composition, we only had a predicate describing that an object $X$ of an abelian category $C$ has projective dimension strictly less than some $n\in\mathbb{N}$, defined as the vanishing of all $\operatorname{Ext}_{i}(X,Y)$ for all $Y$ and all $i\geq n$ (formalized by Professor Joël Riou).
class HasProjectiveDimensionLT (X : C) (n : ℕ) : Prop where mk ::
subsingleton (i : ℕ) (hi : n i) Y : C : Subsingleton (Ext.{max u v} X Y i)
With only this definition, some operations are awkward (for example, adding a number to the projective dimension in the Auslander–Buchsbaum formula). Even if we try to pick out the minimal $n$ such that HasProjectiveDimensionLT X (n + 1), the disadvantages are significant: first, such a number may not exist at all (if $X$ does not have a finite projective resolution); second, finding it with Nat.find complicates operations and requires decidable instances. For these reasons, we want projective dimension as a concrete numeric object.
However, projective dimension cannot be a plain natural number: first, we need to give objects without finite projective resolution the value $\infty$; second, by common convention the projective dimension of the zero object is -$\infty$. We therefore define projective dimension in WithBot ℕ∞:
def projectiveDimension {C : Type u} [Category.{v, u} C] [Abelian C] (X : C) : WithBot ℕ∞ :=
sInf {n : WithBot ℕ∞ | (i : ℕ), n < i HasProjectiveDimensionLT X i}
Although this definition looks somewhat verbose, it behaves according to the usual conventions. (Outside code excerpts we denote projectiveDimension by $\operatorname{projdim}$.)
With this definition, we can prove the equivalence projectiveDimension X < n HasProjectiveDimensionLT X n and many other useful lemmas to make the numeric version convenient in practice. Many lemmas related to HasProjectiveDimensionLT now have corresponding statements for projectiveDimension.
We also prove a lemma specific to the category of $R$-modules that is useful when reducing global dimension to the local case.W
Lemma 2.9.
For a finitely generated module $M$ over Noetherian ring $R$, $\operatorname{projdim}(M)$ is equal to the supremum of $\operatorname{projdim}(M_{\mathfrak{p}})$ over the prime ideals $\mathfrak{p}$ of $R$.
lemma projectiveDimension_eq_iSup_localizedModule_prime
[Small.{v} R] [IsNoetherianRing R] (M : ModuleCat.{v} R) [Module.Finite R M] :
projectiveDimension M = iSup (p : PrimeSpectrum R), projectiveDimension (M.localizedModule p.1.primeCompl)
Proof.
For convenience in formalization, we did not directly prove that $\operatorname{Ext}$ commutes with localization. Noting that the left-hand side being -$\infty$ is equivalent to $M=0$ and similarly for the right-hand side -$\infty$, we instead prove that for each natural number $n$, the statement that the left-hand side is no greater than $n$ is equivalent to the right-hand side being no greater than $n$, by induction on $n$. We construct an exact sequence $0\rightarrow N\rightarrow F\rightarrow M\rightarrow 0$ with $F$ free, localizing at a prime $\mathfrak{p}$ yields an exact sequence $0\rightarrow N_{\mathfrak{p}}\rightarrow F_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}\rightarrow 0$. Using the existing lemma about projective dimension in short exact sequencesW, we reduce the case $n+1$ to the case $n$. We are then left to show that $M$ is projective if and only if each localization at prime $M_{\mathfrak{p}}$ is projective, which is already in mathlib WW. ∎
There is also a version of this lemma considering only maximal ideals. W
Similarly we prove the following lemma.W
Lemma 2.10.
For $S$ a multiplicative set of $R$ and $R$-module $M$, projective dimension of $S^{-1}M$ over $S^{-1}R$ is no greater than projective dimension of $M$ over $R$.
lemma projectiveDimension_le_projectiveDimension_of_isLocalizedModule [Small.{v} R] (S : Submonoid R)
(M : ModuleCat.{v} R) : projectiveDimension (M.localizedModule S) projectiveDimension M
2.6. Global Dimension
Recall that the global dimension of a ring is the supremum of projective dimensions over all $R$-modules. With projective dimension as a numeric object, we implement this definition directly.W
noncomputable def globalDimension : WithBot ℕ∞ :=
iSup (M : ModuleCat.{v} R), projectiveDimension.{v} M
Basic properties and the relation with HasProjectiveDimensionLE are formalized in the same fileW.(Outside code excerpts we denote globalDimension by $\operatorname{gldim}$.)
The supremum over all modules can actually be reduced to the supremum over all finitely generated modules. To prove this we formalized the following form of [16, Lemma 18.2].W
Lemma 2.11.
For a commutative ring $R$ and $n$ a non-negative integer the following are equivalent:
  1. (1)
    $\operatorname{gldim}(R)\leq n$
  2. (2)
    $\operatorname{projdim}(M)\leq n$ for all finitely generated $R$-module $M$
  3. (3)
    $\operatorname{Ext}_{R}^{i}(M,N)$ vanishes for all $R$-module $M,N$ for index $i>n$
  4. (4)
    $\operatorname{injdim}(M)\leq n$ for all $R$-module $M$
Note that in mathlib, projective dimension and injective dimension is already defined via vanishing of $\operatorname{Ext}$; thus, this result might differ slightly from the statement in the usual sense ($(3)\leftrightarrow(4)$ is almost trivial in this setting).
Proof.
$(1)\rightarrow(2)$ and $(3)\rightarrow(1)$ are trivial. For $(2)\rightarrow(3)$, the following proposition suffices.W
Proposition 2.12.
For $R$-module $M$ and $n$ a non-negative integer, if $\operatorname{Ext}_{R}^{n+1}(R/I,M)$ vanish for all ideals $I$ of $R$, then for any $R$-module NN, $\operatorname{Ext}_{R}^{n+1}(N,M)$ vanishes.
Proof.
This result is essentially about the $n$-th syzygy of $M$ in an injective resolution. We prove it by induction on $n$: using dimension-shifting on $M$ we reduce to the case $n=0$, which is just an application of the existing Baer criterionW. ∎
Using Lemma  2.11, we can obtain the following result as a corollary.W
Lemma 2.13 ([8, Theorem 19.1]).
For ring $R$, $\operatorname{gldim}(R)$ is equal to supremum of the projective dimensions of finitely generated modules.
lemma globalDimension_eq_sup_projectiveDimension_finite [Small.{v} R] : globalDimension.{v} R =
iSup (M : ModuleCat.{v} R), iSup (_ : Module.Finite R M), projectiveDimension.{v} M
Combining this with the localization results from the previous subsection, we formalize the following statement.W
Lemma 2.14.
For a noetherian ring $R$, $\operatorname{gldim}(R)$ is equal to the supremum of $\operatorname{gldim}(R_{\mathfrak{p}})$ over the prime ideals $\mathfrak{p}$ of $R$.
lemma globalDimension_eq_iSup_loclization_prime [Small.{v} R] [IsNoetherianRing R] :
globalDimension.{v} R = iSup (p : PrimeSpectrum R), globalDimension.{v} (Localization.AtPrime p.1)
Proof.
Since the left-hand side equals the supremum of projective dimensions of finitely generated modules, by Lemma  2.9, we know that the left is no greater than the right. Conversely, any $R_{\mathfrak{p}}$-module $N$ is already an $R$-module with $N=N_{\mathfrak{p}}$; Lemma  2.10 yields the reverse inequality, completing the proof. ∎
A version for localizing only at maximal ideals is also provided.W
3. Definition of Depth and Related Results
In this section, we formalize the definition of depth as the minimal index of non-vanishing $\operatorname{Ext}$. First, to relate depth to the maximal length of regular sequences contained in a given ideal, we formalize Rees’ theorem (Theorem 3.1). We then establish basic properties of depth, such as the inequalities associated with short exact sequences and the fact that quotienting by a regular element decreases depth by exactly one. Finally, we present our formalizations of two classical results about depth: the Auslander–Buchsbaum theorem in Subsection  3.2 and Ischebeck’s theorem in Subsection  3.3.
Throughout this section, we let $R$ be a Noetherian ring and $M$, NN be $R$-modules.
Theorem 3.1 (Rees, [16, Theorem 28]).
For $M$ a finitely generated $R$-module, and let $I$ be an ideal of $R$ with $IM\neq M$. Fix an integer $n>0$. The following are equivalent.
  1. (1)
    $\operatorname{Ext}^{i}_{R}(N,M)=0$ for all $i<n$, and every finitely generated $R$-module NN with $\mathrm{Supp}(N)\subseteq V(I)$;
  2. (2)
    $\operatorname{Ext}_{R}^{i}(R/I,M)=0$ for all $i<n$;
  3. (3)
    there exists a finite $R$-module NN with $\mathrm{Supp}(N)=V(I)$ such that $\operatorname{Ext}^{i}_{R}(N,M)=0$ for every $i<n$;
  4. (4)
    there exists an $M$-regular sequence $x_{1},\dots,x_{n}$ of length $n$ in $I$.
(The equivalence $(2)\leftrightarrow(4)$ will be a key ingredient in the definition of depth.)
We now present the proof of the main theorem formalized as follows.W
lemma exist_isRegular_tfae [IsNoetherianRing R] (I : Ideal R) [Small.{v} (R I)] (n : ℕ)
(M : ModuleCat.{v} R) (Mntr : Nontrivial M) (Mfin : Module.Finite R M)
(smul_lt : I · (Top.top : Submodule R M) < Top.top) :
[∀ N : ModuleCat.{v} R, (Nontrivial N Module.Finite R N
Module.support R N PrimeSpectrum.zeroLocus I) i < n, Subsingleton (Ext N M i),
i < n, Subsingleton (Ext (ModuleCat.of R (Shrink.{v} (R I))) M i),
N : ModuleCat R, Nontrivial N Module.Finite R N
Module.support R N = PrimeSpectrum.zeroLocus I i < n, Subsingleton (Ext N M i),
rs : List R, rs.length = n (∀ r rs, r I) RingTheory.Sequence.IsRegular M rs ].TFAE
Proof.
The implications $(1)\rightarrow(2)$ and $(2)\rightarrow(3)$ are straightforward. The implications $(3)\rightarrow(4)$ and $(4)\rightarrow(1)$ follow the arguments in [16, Theorem 28]. ∎
The following proposition plays an important role in both of the two main implications and is therefore separated as an independent result for the convenience of the main proof.W
Proposition 3.2.
For finitely generated $M$ and $N$, the $R$-linear maps $N\rightarrow M$ are all zero if and only if there exists an $M$-regular element in the annihilator of $N$.
Proof.
If a prime ideal $\mathfrak{p}\in\mathrm{Ass}(M)$ contains $\mathrm{Ann}(N)$, then the maximal ideal $\mathfrak{p}R_{\mathfrak{p}}$ of $R_{\mathfrak{p}}$ is an associated prime of $M_{\mathfrak{p}}$. We can obtain a non-trivial linear map $N_{\mathfrak{p}}\rightarrow N_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}N_{\mathfrak{p}}\rightarrow R_{\mathfrak{p}}/\mathfrak{m}_{\mathfrak{p}}\rightarrow M_{\mathfrak{p}}$ which yields a non-trivial map $M\rightarrow N$. Applying prime avoidance to $\mathfrak{p}\in\mathrm{Ass}(M)$ then completes the proof. ∎
With Rees’ theorem established, we proceed to the definition of depth.
3.1. Definition of Depth
Rees’ theorem builds a bridge between the homological characterization and the existence of regular sequences. We want a definition of $\operatorname{depth}$ that is meaningful both in commutative algebra and homological terms. We therefore adopt the following definition for the minimal order of non-vanishing $\operatorname{Ext}$ (often called ”grade” in some texts).W
def moduleDepth (N M : ModuleCat.{v} R) : ℕ∞ :=
sSup {n : ℕ∞ | i : ℕ, i < n Subsingleton (Ext N M i)}
Here $R$ is a commutative ring of Type u that is Small.{v} to ensure ModuleCat.{v} R has enough projectives. These assumptions on $R$ are used throughout in this section. (Out side code excerpts, moduleDepth M N will be denoted by $\operatorname{depth}(M,N)$.)
This definition may look unusual but matches the intended idea: it is the minimal index of non-vanishing $\operatorname{Ext}_{R}^{i}(N,M)$, and it equals $\infty$ when no such index exists.
We now specialize to the cases of primary interest. The $\operatorname{depth}$ of a module with respect to an ideal $I$ is defined by taking the module on the left to be $R/I$, formalized as follows:W
def Ideal.depth (I : Ideal R) (M : ModuleCat.{v} R) : ℕ∞ :=
moduleDepth (ModuleCat.of R (Shrink.{v} (R I))) M
For a local ring, $\operatorname{depth}(M)$ denotes the $\operatorname{depth}$ of $M$ with respect to the unique maximal ideal.W
def IsLocalRing.depth [IsLocalRing R] (M : ModuleCat.{v} R) : ℕ∞ :=
(IsLocalRing.maximalIdeal R).depth M
Using Rees’ theorem, we relate the above homological notion to the maximal length of $M$-regular sequences contained in a given ideal. In the statements below, we assume $M$ and $N$ are non-trivial, finitely generated $R$-modules.
Proposition 3.3.
If $I$ an ideal of $R$ with $IM<M$ and $\mathrm{Supp}(N)=V(I)$, then $\operatorname{depth}(N,M)=\operatorname{depth}_{I}(M)$
Proof.
This is a direct application of the implications $(3)\rightarrow(2)\rightarrow(1)$ of Theorem  3.1. ∎
This result relates the general definition of $\operatorname{depth}(M,N)$ with the usual $\operatorname{depth}_{I}(N)$.W
With the same hypothesis on $I$, we have the following relation with length of regular sequences.W
Proposition 3.4.
$\operatorname{depth}(N,M)$ equals the supremum of lengths of $M$-regular sequences contained in $I$.
lemma moduleDepth_eq_sSup_length_regular [IsNoetherianRing R] (I : Ideal R) (N M : ModuleCat.{v} R)
[Module.Finite R M] [Nfin : Module.Finite R N] [Nontrivial M] [Nontrivial N]
(smul_lt : I · (Top.top : Submodule R M) < Top.top)
(hsupp : Module.support R N = PrimeSpectrum.zeroLocus I) :
moduleDepth N M = sSup {(List.length rs : ℕ∞) | (rs : List R)
(_ : RingTheory.Sequence.IsRegular M rs) (_ : r rs, r I) }
Proof.
This is a direct application of the implications $(3)\rightarrow(4)\rightarrow(1)$ of Theorem  3.1. ∎
Returning to the homological definition, we obtain the following inequalities for $\operatorname{depth}$ in a short exact sequence of non-trivial finitely generated $R$-modules.W
Lemma 3.5 (A generalization of [22, 00LX] ).
For a short exact sequence of non-trivial finitely generated $R$ modules $0\rightarrow M^{\prime}\rightarrow M\rightarrow M^{\prime\prime}\rightarrow 0$
  1. (1)
    $\operatorname{depth}(N,M)\geq min\{\operatorname{depth}(N,M^{\prime}),\operatorname{depth}(N,M^{\prime\prime})\}$
  2. (2)
    $\operatorname{depth}(N,M^{\prime\prime})\geq min\{\operatorname{depth}(N,M),\operatorname{depth}(N,M^{\prime})-1\}$
  3. (3)
    $\operatorname{depth}(N,M^{\prime})\geq min\{\operatorname{depth}(N,M),\operatorname{depth}(N,M^{\prime\prime})+1\}$
Proof.
These follow readily from the long exact sequence:
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(N,M^{\prime})\rightarrow\operatorname{Ext}_{R}^{i}(N,M)\rightarrow\operatorname{Ext}_{R}^{i}(N,M^{\prime\prime})\rightarrow\operatorname{Ext}_{R}^{i+1}(N,M^{\prime})\rightarrow\operatorname{Ext}_{R}^{i+1}(N,M)\rightarrow\cdots$
The corresponding ‘dual’ statements about $\operatorname{depth}(\cdot,M)$ in short exact sequences are proved similarly using the contravariant long exact sequence.W However, for Ideal.depth and IsLocalRing.depth, the dual version is rarely used.
We also formalized the behavior of depth under quotient by a regular element.W
Theorem 3.6 (A generalization of [22, 090R]).
For an $M$-regular element $x\in\mathrm{Ann}(N)$, we have
$\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$
lemma moduleDepth_quotSMulTop_succ_eq_moduleDepth
(N M : ModuleCat.{v} R) (x : R) (reg : IsSMulRegular M x) (mem : x Module.annihilator R N) :
moduleDepth N (ModuleCat.of R (QuotSMulTop x M)) + 1 = moduleDepth N M
Here QuotSMulTop x M denotes the quotient module $M/xM$.
Proof.
From the short exact sequence $0\rightarrow M\xrightarrow{\cdot x}M\rightarrow M/xM\rightarrow 0$, we obtain the long exact sequence
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i}(N,M)\rightarrow\operatorname{Ext}_{R}^{i}(N,M/xM)\rightarrow\operatorname{Ext}_{R}^{i+1}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i+1}(N,M)\rightarrow\cdots$
Since $x\in Ann(N)$, the maps $\operatorname{Ext}_{R}^{i}(N,M)\xrightarrow{\cdot x}\operatorname{Ext}_{R}^{i}(N,M)$ are zero for all $i$, so $\operatorname{Ext}_{R}^{i}(N,M/xM)$ vanish if and only if both $\operatorname{Ext}_{R}^{i}(N,M)$ and $\operatorname{Ext}_{R}^{i+1}(N,M)$ vanish. The theorem follows from this observation. ∎
As a corollary, there is an analogous statement for quotient by a (weakly) regular sequence. W
Corollary 3.7.
For $(r_{1},\cdots,r_{k})$ an $M$-regular sequence contained in $\mathrm{Ann}(N)$, then
$\operatorname{depth}(N,M)+k=\operatorname{depth}(N,M/(r_{1},\cdots,r_{k})M)$
Proof.
This result follows easily by induction on the length of the sequence. ∎
3.2. The Auslander–Buchsbaum Theorem
In this section, we prove the following theorem first introduced in [2].W
Theorem 3.8 (Auslander–Buchsbaum).
Let $(R,\mathfrak{m})$ be a Noetherian local ring and let $M$ be a non-trivial finitely generated $R$-module. If projective dimension of $M$ is finite, then
$\operatorname{projdim}(M)+\operatorname{depth}(M)=\operatorname{depth}(R)$
Proof.
We first establish a preliminary result.W
Proposition 3.9.
For a finitely generated free module $M$ over $R$, for any $R$-module NN, $\operatorname{Ext}_{R}^{i}(N,M)$ vanish if and only if $\operatorname{Ext}_{R}^{i}(N,R)$ vanish.
Proof.
This can be deduced from the fact that $\operatorname{Ext}$ commutes with finite bi-productsW. ∎
As a corollary, $\operatorname{depth}(M)=\operatorname{depth}(R)$ for non-trivial finitely generated free module $M$.W
Returning to the proof of the main theorem, the case $\operatorname{projdim}(M)=0$ is immediate since then $M$ is projective and hence free, the equality follows from the above results.
The case $\operatorname{projdim}(M)=1$ is the core, so we prove it separately, following the proof in [14, Theorem 19.1].
lemma AuslanderBuchsbaum_one [IsNoetherianRing R] [IsLocalRing R] (M : ModuleCat.{v} R) [Nontrivial M]
[Module.Finite R M] (le1 : HasProjectiveDimensionLE M 1) (nle0 : ¬ HasProjectiveDimensionLE M 0) :
1 + IsLocalRing.depth M = IsLocalRing.depth.{v} (ModuleCat.of.{v} R (Shrink.{v} R))
Pick a $k$ basis of $M/\mathfrak{m}M$ and lift it back to $M$, obtaining a map $f:R^{n}\rightarrow M$ whose induced map $(R/\mathfrak{m})^{n}\rightarrow M/\mathfrak{m}M$ is an isomorphism; hence $\mathrm{ker}(f)\subseteq\mathfrak{m}R^{n}$. Since $\mathrm{ker}(f)$ has projective dimension 0, it is projective thus free.
In ordinary proofs, we could view $f$ as a matrix with all elements contained in $\mathfrak{m}$. However, during formalization, viewing a map between free modules as a matrix and operating with it would be complicated, so instead we set up the following two propositions to help.WW
Proposition 3.10.
For $M$ a finitely generated free module over $R$, if the range of $f:M\rightarrow N$ is contained in $IN$ where $I$ is an ideal of $R$, then $f\in I\cdot(M\rightarrow N)$
Proposition 3.11.
For $R$-modules $L,M,N$, if $f:M\rightarrow N$ is contained in $\mathrm{Ann}(L)\cdot(M\rightarrow N)$, then the map $\operatorname{Ext}_{R}^{i}(L,M)\rightarrow\operatorname{Ext}_{R}^{i}(L,N)$ induced by $f$ is zero.
Proof.
One verifies this lemma easily by reducing to the case $f=a\cdot g$ with $a\in\mathrm{Ann}(L)$. ∎
Then we know that $i:\mathrm{ker}(f)\hookrightarrow R^{n}$ is contained in $\mathfrak{m}\cdot(\mathrm{ker}(f)\rightarrow R^{n})$, so the second lemma shows that the map $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})$ induced by $i:\mathrm{ker}(f)\hookrightarrow R^{n}$ is zero.
Now consider the long exact sequence induced by $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$:
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})\rightarrow\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i+1}(k,R^{n})\rightarrow\cdots$
The first and fourth maps are zero by the discussion above. Moreover, vanishing of $\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))$ and of $\operatorname{Ext}_{R}^{i}(k,R^{n})$ are each equivalent to vanishing of $\operatorname{Ext}_{R}^{i}(k,R)$. From the long exact sequence, we deduce that $\operatorname{Ext}_{R}^{i}(k,M)$ vanishes if and only if both $\operatorname{Ext}_{R}^{i}(k,R)$ and $\operatorname{Ext}_{R}^{i+1}(k,R)$ vanish. By the homological characterization of depth, this yields $1+\operatorname{depth}(M)=\operatorname{depth}(R)$.
Finally, we prove the theorem by induction on $\operatorname{projdim}(M)$. Assume the statement holds for projective dimension $n\geq 1$, and let $\operatorname{projdim}(M)=n+1$. Take a surjection $f:R^{m}\rightarrow M$; then $\mathrm{ker}(f)$ has projective dimension $n$. By induction hypothesis, $n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$. Since $\operatorname{depth}(R^{m})=\operatorname{depth}(R)\geq n>0$, so there is no non-trivial $k\rightarrow\mathrm{ker}(f)\hookrightarrow R^{m}$, thus $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$ vanish.
Consider the long exact sequence associated to $0\rightarrow\mathrm{ker}(f)\rightarrow R^{n}\rightarrow M\rightarrow 0$:
$\cdots\rightarrow\operatorname{Ext}_{R}^{i}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i}(k,R^{n})\rightarrow\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))\rightarrow\operatorname{Ext}_{R}^{i+1}(k,R^{n})\rightarrow\cdots$
For all $i<\operatorname{depth}(\mathrm{ker}(f))<\operatorname{depth}(R)$, the map $\operatorname{Ext}_{R}^{i}(k,M)\rightarrow\operatorname{Ext}_{R}^{i+1}(k,\mathrm{ker}(f))$ is an isomorphism. Together with the vanishing of $\operatorname{Ext}_{R}^{0}(k,\mathrm{ker}(f))$, we obtain $\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(M)+1$. Hence $n+1+\operatorname{depth}(M)=n+\operatorname{depth}(\mathrm{ker}(f))=\operatorname{depth}(R)$, which completes the induction and the proof. ∎
3.3. The Ischebeck Theorem
With the general definition moduleDepth, we can state the Ischebeck’s theorem (first introduced in [12]) as follows.W
Theorem 3.12 (Ischbeck).
For Noetherian local ring $R$ and non-trivial finitely generated module $M$ and $N$, $\operatorname{depth}(M,N)\geq\operatorname{depth}(M)-\dim(N)$
theorem moduleDepth_ge_depth_sub_dim [IsNoetherianRing R] [IsLocalRing R] (M N : ModuleCat.{v} R)
[Module.Finite R M] [Nfin : Module.Finite R N] [Nontrivial M] [Nntr : Nontrivial N] [Small.{v} R] :
moduleDepth N M IsLocalRing.depth M -
(Module.supportDim R N).unbot (Module.supportDim_ne_bot_of_nontrivial R N)
Proof.
Since the case $\dim(N)=\infty$ is not possible, we prove the statement by strong induction on $\dim(N)$. If $\dim(N)=0$, then $\mathrm{Supp}(N)=V(\mathfrak{m})$, and hence $\operatorname{depth}(N,M)=\operatorname{depth}(M)$ by Proposition  3.3.
Now assume the statement holds for modules of dimension no greater than $n$, and let $\dim(N)=n+1$. In the development of associated primes, one has an induction principle for finitely generated modules over a Noetherian ringW. It suffices to check the property in the following cases: $N=0$ (irrelevant here), $N\cong R/\mathfrak{p}$ for some prime ideal $\mathfrak{p}$, and that the property is preserved under extension by short exact sequences. Preservation under extension by short exact sequences follows easily from Lemma  3.5 and a straightforward calculation. Care must be taken to apply the induction hypothesis appropriately to $N_{1}$ in the two subcases when $\dim(N_{1})\leq n$ and when $\dim(N_{1})=n+1$, because in this process the cited induction principle only provides hypotheses for modules of dimension $n+1$. The hypotheses in the former case come from the strong induction on dimension (and similarly for $N_{3}$). Thus the remaining and essential case is $N\cong R/\mathfrak{p}$, which is handled by the same argument as in [16, Lemma 15.2]. ∎
Using this theorem, we obtain a useful bound on $\operatorname{depth}$.W
Corollary 3.13 ([16, Theorem 29]).
For finitely generated $R$-module $M$ and any $\mathfrak{p}\in\mathrm{Ass}(M)$, $\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})$
theorem depth_le_ringKrullDim_associatedPrime [IsNoetherianRing R] [IsLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] [Nontrivial M] (P : Ideal R) (ass : P associatedPrimes R M) :
IsLocalRing.depth M (ringKrullDim (R P)).unbot (quotient_prime_ringKrullDim_ne_bot ass.1)
Proof.
Since $\operatorname{depth}(M,R/\mathfrak{p})=0$, by Theorem  3.12, we have $\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})$. ∎
Moreover, from this we deduce the well-known inequalities $\operatorname{depth}(M)\leq\dim(M)\leq\dim(R)<\infty$.W
4. Cohen–Macaulay Modules and Rings
In this section, we first define Cohen–Macaulay modules over local rings. We then prove that passing to the quotient by a regular element preserves the Cohen–Macaulay property, and that localization of a Cohen–Macaulay module remains Cohen–Macaulay. Similarly, we establish corresponding results for Cohen–Macaulay rings. In Subsection  4.1, we prove the unmixedness theorem for Cohen–Macaulay rings; as a byproduct, we obtain that a Cohen–Macaulay local ring is catenary. Since a polynomial ring over a Cohen–Macaulay ring is again Cohen–Macaulay (proved in Subsection  4.2), it will be straightforward to deduce that a Cohen–Macaulay ring is universally catenary once the preliminary constructions are in place.
With a well-developed definition of depth, we define a Cohen–Macaulay module (over local ring) following [16, 16.A].W
Definition 4.1.
An $R$-module $M$ is a Cohen–Macaulay module if $M=0$ or $\operatorname{depth}(M)=\dim(M)$.
class ModuleCat.IsCohenMacaulay [IsLocalRing R] [Small.{v} R] (M : ModuleCat.{v} R) : Prop where
depth_eq_dim : Subsingleton M Module.supportDim R M = IsLocalRing.depth M
Corollary  3.13 yields the following useful lemma, formalized as two separate statements.W W
Lemma 4.2 ([16, Theorem 30(i)]).
For $M$ a finitely generated Cohen–Macaulay module over Noetherian ring $R$, for any $\mathfrak{p}\in\mathrm{Ass}(M)$, $\dim(M)=\operatorname{depth}(M)=\dim(R/\mathfrak{p})$. Consequently $\mathfrak{p}$ is minimal prime over $\mathrm{Ann}(M)$.
Proof.
Since $\mathrm{Ann}(M)\subseteq\mathfrak{p}$, by Corollary  3.13, we have $\dim(M)=\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})\leq\dim(R/\mathrm{Ann}(M))=\dim(M)$, so every inequality in the chain must be an equality. Hence $\dim(M)=\operatorname{depth}(M)=\dim(R/\mathfrak{p})$, forcing $\mathfrak{p}$ to be a minimal prime over $\mathrm{Ann}(M)$. ∎
Combining this with Lemma  2.2, we find that the associated primes of a Cohen–Macaulay module are exactly the minimal primes over $\mathrm{Ann}(M)$.W
lemma associated_prime_eq_minimalPrimes_isCohenMacaulay
(M : ModuleCat.{v} R) [M.IsCohenMacaulay] [Module.Finite R M] [Nontrivial M] :
associatedPrimes R M = (Module.annihilator R M).minimalPrimes
For a finitely generated $R$-module $M$ with $R$ local, quotienting by an $M$-regular element in the maximal ideal reduces both the depth and the Krull dimension of $M$ by exactly one; therefore, whether $M$ is Cohen–Macaulay is preserved under such a quotient.W
Lemma 4.3.
For $M$ a finitely generated module over $R$ and $x$ an $M$-regular element, then $M$ is Cohen–Macaulay if and only if $M/xM$ is Cohen–Macaulay.
Proof.
This is a direct consequence of Theorem  3.6 and Lemma  2.3. ∎
A similar result holds for quotienting by a regular sequence, which would be used later in subsection  4.1.W
Lemma 4.4 ([16, Theorem 30(ii)]).
For $M$ a finitely generated module over $R$ and $(r_{1},\cdots,r_{k})$ an $M$-regular sequence, then $M$ is Cohen–Macaulay if and only if $M/(r_{1},\cdots,r_{k})M$ is Cohen–Macaulay.
lemma quotient_regular_isCohenMacaulay_iff_isCohenMacaulay
(M : ModuleCat.{v} R) [Module.Finite R M] (rs : List R) (reg : IsRegular M rs) :
M.IsCohenMacaulay (ModuleCat.of R (M Ideal.ofList rs · (Top.top : Submodule R M))).IsCohenMacaulay
Proof.
This can be proved easily by induction on the length of the sequence. ∎
We then formalized that localization of a Cohen–Macaulay module at a prime ideal is again Cohen–Macaulay.
Lemma 4.5 ([16, Theorem 30(iii)]).
For $M$ a finitely generated Cohen–Macaulay module over $R$ and $\mathfrak{p}$ a prime ideal of $R$, then $M_{\mathfrak{p}}$ is Cohen–Macaulay over $R_{\mathfrak{p}}$. Furthermore, $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$.
This result is formalized as the following two separate statements.WW
lemma isLocalize_at_prime_depth_eq_of_isCohenMacaulay (p : Ideal R) [p.IsPrime] [IsLocalRing Rₚ]
[Module.Finite R M] [M.IsCohenMacaulay] : Mₚ.IsCohenMacaulay
lemma isLocalize_at_prime_dim_eq_prime_depth_of_isCohenMacaulay [Module.Finite R M] [M.IsCohenMacaulay]
[ntr : Nontrivial Mₚ] : Module.supportDim R M = p.depth M
Proof.
To prove this result, we first set up a lemma.W
Proposition 4.6.
For $M$ a finitely generated module over $R$ and $\mathfrak{p}$ a prime ideal of $R$, $\operatorname{depth}_{\mathfrak{p}}(M)\leq\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$.
Proof.
Here we use the regular sequence characterization of depth. As mentioned in Subsection  2.4, the image of an $M$-regular sequence contained in $\mathfrak{p}$ under the map $R\to R_{\mathfrak{p}}$ is $M_{\mathfrak{p}}$-regular, yielding the above inequality. ∎
Returning to the main proof, since we know $\operatorname{depth}(M_{\mathfrak{p}})\leq\dim(M_{\mathfrak{p}})$ from Subsection  3.3, to prove $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$, it remains to show $\operatorname{depth}_{\mathfrak{p}}(M)=\dim(M_{\mathfrak{p}})$.
From subsection  3.3, we know that $\operatorname{depth}_{\mathfrak{p}}(M)\leq\operatorname{depth}(M_{\mathfrak{p}})$ is finite, so we proceed by induction on $\operatorname{depth}_{\mathfrak{p}}(M)$ following the proof in [16, Theorem 30(iii)]. This completes the proof that the localization at a prime of a Cohen–Macaulay module is Cohen–Macaulay, together with the equality $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$. ∎
The formalization of a Cohen–Macaulay local ring is done more directly rather than by extending the module notion.W
Definition 4.7.
A local ring $R$ is a Cohen–Macaulay local ring if $\operatorname{depth}(R)=\dim(R)$.
class IsCohenMacaulayLocalRing : Prop extends IsLocalRing R where
depth_eq_dim : ringKrullDim R = IsLocalRing.depth (ModuleCat.of R R)
This is equivalent to (ModuleCat.of R R).IsCohenMacaulay, but avoids passing through dimension of modules.W
From the result that localization of a Cohen–Macaulay module at a prime is Cohen–Macaulay, we obtain the corresponding version for Cohen–Macaulay local rings as a corollary.W
Corollary 4.8.
For a Cohen–Macaulay local ring $R$ and $\mathfrak{p}$ a prime ideal of $R$, then $R_{\mathfrak{p}}$ is Cohen–Macaulay.
lemma isCohenMacaulayLocalRing_localization_atPrime [IsCohenMacaulayLocalRing R] [IsNoetherianRing R]
(p : Ideal R) [p.IsPrime] (R : Type*) [CommRing Rₚ] [Algebra R Rₚ] [IsLocalization.AtPrime R p] :
IsCohenMacaulayLocalRing R
We also establish the corresponding results of Lemma  4.3 and Lemma  4.4 for Cohen–Macaulay local rings.WW
Corollary 4.9.
For a local ring $R$ and an $R$-regular element $x$, $R$ is Cohen–Macaulay if and only if $R/(x)$ is Cohen–Macaulay.
Corollary 4.10.
For a local ring $R$ and an $R$-regular sequence $(r_{1},\cdots,r_{k})$, $R$ is Cohen–Macaulay if and only if $R/(r_{1},\cdots,r_{k})$ is Cohen–Macaulay.
With Definition  4.7, we can define general (non-local) Cohen–Macaulay rings.W
Definition 4.11.
A ring $R$ is a Cohen–Macaulay ring if $R_{\mathfrak{p}}$ is a Cohen–Macaulay local ring for every prime ideal $\mathfrak{p}$.
class IsCohenMacaulayRing : Prop where
CM_localize : p : Ideal R, (_ : p.IsPrime), IsCohenMacaulayLocalRing (Localization.AtPrime p)
Using the above localization result for Cohen–Macaulay local rings, we find that the Cohen–Macaulay property need only be verified at maximal ideals.W
Lemma 4.12.
A Noetherian ring $R$ is Cohen–Macaulay if and only if $R_{\mathfrak{m}}$ is Cohen–Macaulay local ring for every maximal ideal $\mathfrak{m}$.
lemma isCohenMacaulayRing_iff [IsNoetherianRing R] : IsCohenMacaulayRing R
m : Ideal R, (_ : m.IsMaximal), IsCohenMacaulayLocalRing (Localization.AtPrime m)
As a corollary, a Cohen–Macaulay local ring is a Cohen–Macaulay ringW and a local ring that is a Cohen–Macaulay ring is a Cohen–Macaulay local ring W.
4.1. Unmixedness Theorem for Cohen–Macaulay Rings
The unmixedness theorem for Cohen–Macaulay rings was first stated in the graded case by Macaulay in [13]. Here we follow the more modern formulation originating from [17].
We first prove the following Theorem.
Theorem 4.13 ([16, Theorem 31(iii)]).
Let $R$ be a Cohen–Macaulay local ring and let $a_{1},\dots,a_{r}$ be a sequence in the unique maximal ideal $\mathfrak{m}$. The following are equivalent.
  1. (1)
    $a_{1},\cdots,a_{r}$ is $R$-regular
  2. (2)
    $0pt(a_{1},\cdots,a_{i})=i$, for all $1\leq i\leq r$
  3. (3)
    $0pt(a_{1},\cdots,a_{r})=r$
  4. (4)
    $a_{1},\cdots,a_{r}$ can be extended to a system of parameters.
We do not state the theorem as a single List.TFAE, but instead provide step-by-step lemmas, since some parts do not essentially use the Cohen–Macaulay or even the local hypothesis.W W W In the following lemmas, we assume $R$ is Noetherian.
Proposition 4.14 (Theorem  4.13 $(1)\rightarrow(2)$).
For ring $R$ and a weakly regular sequence $(r_{1},\cdots,r_{k})$ generating a proper ideal, then $0pt(r_{1},\cdots,r_{k})=k$ .
lemma Ideal.ofList_height_eq_length_of_isWeaklyRegular (rs : List R) (reg : IsWeaklyRegular R rs)
(h : Ideal.ofList rs Top.top) : (Ideal.ofList rs).height = rs.length
Proof.
The result follows easily by induction on the length of the sequence using Krull’s height theorem and the fact that a regular element is not contained in any minimal prime. ∎
Proposition 4.15 (Theorem  4.13 $(3)\rightarrow(4)$).
For local ring $R$ and a sequence $(r_{1},\cdots,r_{k})$, if $0pt(r_{1},\cdots,r_{k})=k$ then it can be extended to a system of parameters.
lemma maximalIdeal_mem_ofList_append_minimalPrimes_of_ofList_height_eq_length [IsLocalRing R]
(rs : List R) (mem : r rs, r maximalIdeal R) (ht : (Ideal.ofList rs).height = rs.length) :
rs : List R, maximalIdeal R (Ideal.ofList (rs ++ rs’)).minimalPrimes
rs.length + rs’.length = ringKrullDim R
Proof.
This result follows by adding elements to the tail inductively while ensuring that the height of the ideal generated by the sequence remains equal to the length of the sequence. ∎
Proposition 4.16 (Preparation for Theorem  4.13 $(4)\rightarrow(1)$).
For a Cohen–Macaulay local ring $R$, a system of parameters $(r_{1},\cdots,r_{k})$ form a regular sequence.
lemma isRegular_of_maximalIdeal_mem_ofList_minimalPrimes [IsCohenMacaulayLocalRing R]
(rs : List R) (mem : maximalIdeal R (Ideal.ofList rs).minimalPrimes)
(dim : rs.length = ringKrullDim R) : IsRegular R rs
Proof.
If the first $k$ elements $(r_{1},\cdots,r_{k})$ form a regular sequence, then by Corollary  4.10, $R/(r_{1},\cdots,r_{k})$ is Cohen–Macaulay. As all associated primes of $R/(r_{1},\cdots,r_{k})$ are minimal by Lemma  4.2, the fact that $r_{k+1}$ is not contained in any minimal prime of $R/(r_{1},\cdots,r_{k})$ implies it is a regular element. ∎
Propositions  4.15 and  4.16 together imply the following result.W
Proposition 4.17 (Theorem  4.13 $(4)\rightarrow(1)$).
For local ring $R$ and a sequence $(r_{1},\cdots,r_{k})$, if $0pt(r_{1},\cdots,r_{k})=k$ then it is a regular sequence.
lemma isRegular_of_ofList_height_eq_length_of_isCohenMacaulayLocalRing [IsCohenMacaulayLocalRing R]
(rs : List R) (mem : r rs, r maximalIdeal R) (ht : (Ideal.ofList rs).height = rs.length) :
IsRegular R rs
A by-product of Theorem  4.13 is the following lemma.W
Lemma 4.18 ([16, Theorem 31(i), Part1]).
For $R$ a Cohen–Macaulay local ring and $I$ a proper ideal of it, then $\operatorname{depth}_{I}(R)=0pt(I)$.
lemma Ideal.depth_eq_height [IsCohenMacaulayLocalRing R] (I : Ideal R) (netop : I Top.top) :
I.depth (ModuleCat.of R R) = I.height
Proof.
By Theorem  4.13, any regular sequence in $I$ has length at most $0pt(I)$, so the left-hand side is not greater than the right-hand side. Then as we can find $J\subseteq I$ such that $0pt(J)=0pt(I)$ and $J$ can be generated by $0pt(I)$ elements, Theorem  4.13 shows the generators form a regular sequence, then the right-hand side is not greater than the left-hand side. ∎
We then establish the following theorem which implies that a Cohen–Macaulay local ring is catenary. W
Theorem 4.19 ([16, Theorem 31(i), Part2]).
For $R$ a Cohen–Macaulay local ring and $I$ a proper ideal, we have $0pt(I)+\dim(R/I)=\dim(R)$.
lemma Ideal.height_add_ringKrullDim_quotient_eq_ringKrullDim [IsCohenMacaulayLocalRing R]
(I : Ideal R) (netop : I Top.top) : I.height + ringKrullDim (R I) = ringKrullDim R
Proof.
Since minimal primes over $I$ is finite, by simple manipulations of iSup and iInf we can reduce to the case where $I$ is a prime ideal.W The formalization then follows the proof in [16, Theorem 31(i), Part 2]. ∎
From this result and the fact that localization of a Cohen–Macaulay ring is still Cohen–Macaulay, for prime ideals $\mathfrak{q}\subseteq\mathfrak{p}$, we have $\dim(R_{\mathfrak{p}})=0pt(\mathfrak{q}R_{\mathfrak{p}})+\dim(R_{\mathfrak{p}}/\mathfrak{q}R_{\mathfrak{p}})$, i.e. $0pt(\mathfrak{p})-0pt(\mathfrak{q})=0pt(\mathfrak{p}/\mathfrak{q})$. This implies that a Cohen–Macaulay local ring is catenary. However, since notions of catenary order and catenary rings are not yet in mathlib, we do not develop this further here.
Returning to the main theorem, we first define the notion of an unmixed ideal.W
Definition 4.20.
For $I$ an ideal of a Noetherian ring $R$, II is unmixed if for any $\mathfrak{p}\in\mathrm{Ass}(R/I)$, $0pt(\mathfrak{p})=0pt(I)$.
class Ideal.IsUnmixed (I : Ideal R) : Prop where
height_eq : {p : Ideal R}, p associatedPrimes R (R I) p.height = I.height
In particular, if $I=(r_{1},\cdots,r_{k})$ with $0pt(I)=k$, then $I$ is unmixed if and only if $R/I$ has no embedded primes.
The unmixedness theorem can then be stated as the follows.
Theorem 4.21.
A Noetherian ring $R$ is Cohen–Macaulay if and only if for any ideal $I=(r_{1},\cdots,r_{k})$ with $0pt(I)=k$, II is unmixed.
theorem isCohenMacaulayRing_iff_unmixed : IsCohenMacaulayRing R
(l : List R), (Ideal.ofList l).height = l.length (Ideal.ofList l).IsUnmixed
Proof.
With the results earlier in this section, the formalization of the unmixedness theorem follows easily from the proof in [16, Theorem 32].W
4.2. Universally Catenary
Using the results proven in Subsection  4.1, to prove that Cohen–Macaulay rings are universally catenary, it remains only to show that a polynomial ring over a Cohen–Macaulay ring is also Cohen–Macaulay.W
theorem Polynomial.isCM_of_isCM [IsNoetherianRing R] [IsCohenMacaulayRing R] : IsCohenMacaulayRing R[X]
Assume $R$ is Cohen–Macaulay. For any prime ideal $\mathfrak{p}$ of $R[X]$, set $\mathfrak{q}=\mathfrak{p}\cap R$. Note that $R[X]_{\mathfrak{p}}$ is the localization of $R_{\mathfrak{q}}[X]$ at $\mathfrak{p}R_{\mathfrak{q}}[X]$, so the problem reduces to the following case: if $R$ is a Cohen–Macaulay local ring and $\mathfrak{p}$ is a prime ideal of $R[X]$ with $\mathfrak{p}\cap R$ equal to the maximal ideal of $R$, then $R[X]_{\mathfrak{p}}$ is a Cohen–Macaulay local ring.W
lemma Polynomial.localization_at_comap_maximal_isCM_isCM [IsNoetherianRing R] [IsCohenMacaulayLocalRing R]
(p : Ideal R[X]) [p.IsPrime] (max : p.comap C = maximalIdeal R) :
IsCohenMacaulayLocalRing (Localization.AtPrime p)
Using the preservation of (weakly) regular sequences under flat base change discussed in Subsection  2.4, the proof follows the argument in [16, Theorem 33].
Although the notions of catenary order and catenary rings are not yet available in mathlib, once they are added, it will be straightforward to combine the above results to conclude that Cohen–Macaulay rings are universally catenary.
5. Regular Local Ring
In this section, we first introduce our formalization of the definition of regular local rings and prove that they are domains. Then we prove the two directions of the Auslander–Buchsbaum–Serre criterion in the next two subsections. We first prove that the global dimension of a regular local ring equals its Krull dimension by showing that a maximal Cohen–Macaulay module over a regular local ring is free. Then we prove the converse: if the unique maximal ideal of a Noetherian local ring has finite projective dimension, then it is generated by a regular sequence; this completes the converse direction and can be viewed as a weakened form of the Ferrand–Vasconcelos theorem. Finally, we establish some corollaries of the Auslander–Buchsbaum–Serre criterion in the last subsection, including that regularity of a ring can be checked at maximal ideals only, and Hilbert’s Syzygy Theorem.
We define a regular local ring simply as a Noetherian local ring whose maximal ideal has span rank (minimal number of generators) equal to its Krull dimension.W
class IsRegularLocalRing : Prop extends IsLocalRing R, IsNoetherianRing R where
reg : (maximalIdeal R).spanFinrank = ringKrullDim R
Using the fact that the span rank of the maximal ideal equals the dimension of the cotangent space over the residue fieldW, we establish the equivalence of this definition with the cotangent-space formulationW.
lemma IsRegularLocalRing.iff_finrank_cotangentSpace [IsLocalRing R] [IsNoetherianRing R] :
IsRegularLocalRing R Module.finrank (ResidueField R) (CotangentSpace R) = ringKrullDim R
5.1. Regular Local Ring is Domain
To prove that a regular local ring is a domain and that minimal generators of the maximal ideal form a regular sequence, we establish the following lemma.W
Lemma 5.1.
Let $(R,\mathfrak{m},k)$ be a regular local ring of dimension $d$, and let $a_{1},\cdots,a_{k}\in\mathfrak{m}$. Then the followings are equivalent.
  1. (1)
    $a_{1},\cdots,a_{k}$ can be extended to a regular system of parameters of $R$.
  2. (2)
    The images $\overline{a_{1}},\cdots,\overline{a_{k}}\in\mathfrak{m}/\mathfrak{m}^{2}$ are $k$-linearly independent.
  3. (3)
    $R/(a_{1},\cdots,a_{k})$ is a regular local ring of dimension dkd-k.
lemma quotient_isRegularLocalRing_tfae
[IsRegularLocalRing R] (S : Finset R) (sub : (S : Set R) maximalIdeal R) :
[∃ (T : Finset R), S T T.card = ringKrullDim R Ideal.span T = maximalIdeal R,
LinearIndependent (ResidueField R) (((maximalIdeal R).toCotangent).comp (Set.inclusion sub)),
IsRegularLocalRing (R Ideal.span (S : Set R))
(ringKrullDim (R Ideal.span (S : Set R)) + S.card = ringKrullDim R) ].TFAE
Proof.
For $(1)\rightarrow(2)$: if $a_{1},\cdots,a_{k}$ extend to $a_{1},\cdots,a_{d}$ generating $\mathfrak{m}$, their images in the cotangent space are linearly independent, so $(2)$ holds.
For $(2)\rightarrow(3)$: let $\tilde{\mathfrak{m}}$ be the maximal ideal of $R/(a_{1},\cdots,a_{k})$. Then $\tilde{\mathfrak{m}}/\tilde{\mathfrak{m}}^{2}$ is isomorphic to $\mathfrak{m}/\mathfrak{m}^{2}/(\overline{a_{1}},\cdots,\overline{a_{k}})$, which has dimension dkd-k. Since the dimension of $R/(a_{1},\dots,a_{k})$ drops by at most $k$ from that of $R$, equality follows.
For $(3)\rightarrow(1)$: pick generators $\tilde{a_{k+1}},\cdots,\tilde{a_{d}}$ of $\tilde{\mathfrak{m}}$ and lift them to $R$; together with $a_{1},\cdots,a_{k}$, these elements generate $\mathfrak{m}$, hence form a regular system of parameters. ∎
In particular, this theorem yields the following result for the quotient by a single element.W
Corollary 5.2.
For regular local ring $R$, if $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$, then $R/(x)$ is a regular local ring of dimension $\dim(R)-1$.
We then prove that a regular local ring $(R,\mathfrak{m},k)$ is a domain by induction on the dimension.W
Theorem 5.3.
A regular local ring is a domain.
theorem isDomain_of_isRegularLocalRing [IsRegularLocalRing R] : IsDomain R
Proof.
Proof by induction on $\dim(R)$. Pick $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ that is not in any minimal prime of $R$. The induction hypothesis, along with Corollary  5.2 shows that $R/(x)$ is a domain. From the property of $x$, we know that $(x)$ is a prime ideal that is not minimal. The rest of the proof follows the arguments in [5, Proposition 2.2.3]. ∎
For $R$ regular local ring of dimension $d$ and a sequence $rs$ of length $d$ generating its maximal ideal, any quotient of $R$ by a subsequence of $rs$ is still regular and hence a domain by the result above; therefore, $rs$ is an $R$-regular sequence.W
theorem isRegular_of_span_eq_maximalIdeal [IsRegularLocalRing R] (rs : List R)
(span : Ideal.ofList rs = maximalIdeal R) (len : rs.length = ringKrullDim R) : IsRegular R rs
As a corollary, a regular local ring is Cohen–Macaulay.W
5.2. Global Dimension of Regular Local Ring
In this section, we prove that the global dimension of a regular local ring equals its Krull dimension.W
Theorem 5.4.
For regular local ring $R$, $\operatorname{gldim}(R)=\dim(R)$.
theorem IsRegularLocalRing.globalDimension_eq_ringKrullDim [Small.{v} R] [IsRegularLocalRing R] :
globalDimension.{v} R = ringKrullDim R
Currently the definition of globalDimension depends on the universe of the category, further discussion appears in subsection  6.1.
There is a well-known proof of this theorem using the following two facts for a Noetherian local ring $(R,\mathfrak{m},k)$.
  • $\operatorname{gldim}(R)=\operatorname{projdim}(k)$ [16, Theorem 41]
  • For any finitely generated $R$-module $M$ and an $M$-regular element $x$, we have
    $\operatorname{projdim}(M/xM)=\operatorname{projdim}(M)+1$[16, Lemma 18.6]
Considering $k$ as the quotient of $R$ by a regular sequence completes the proof. However, the $\operatorname{Tor}$ functor and related infrastructure were still incomplete in mathlib when this work was finished, so we could not prove the first fact with the current setup. Instead, we prove that maximal Cohen–Macaulay modules over a regular local ring are free and use that to proceed.
5.2.1. Maximal Cohen–Macaulay Module
We first state a lemma used both here and in the next subsection.W
Lemma 5.5 ([22, 00NS]).
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring and $M$ a finitely generated $R$-module. Let $x\in\mathfrak{m}$ be $M$-regular. If $M/xM$ is free over $R/(x)$, then $M$ is free over $R$.
lemma free_of_quotSMulTop_free [IsLocalRing R] [IsNoetherianRing R] (M : Type*) [AddCommGroup M]
[Module R M] [Module.Finite R M] {x : R} (mem : x maximalIdeal R) (reg : IsSMulRegular M x)
(free : Module.Free (R Ideal.span {x}) (QuotSMulTop x M)) : Module.Free R M
Proof.
(Sketch) Pick the basis of $M/xM$ over $R/(x)$ and lift it back to $M$, from this we can obtain a surjection $f:R^{n}\rightarrow M$ with $\mathrm{ker}(f)\subseteq xR^{n}$. As $x$ is $M$-regular, $\mathrm{ker}(f)=x\mathrm{ker}(f)$. By Nakayama lemma, this implies $f$ is isomorphism. For further details, refer to the argument in [22, 00NS]. ∎
Returning to the main thread, we define a maximal Cohen–Macaulay module over $R$ as a module whose depth equals $\dim(R)$.W
Definition 5.6.
A module $M$ over local ring $R$ is a maximal Cohen–Macaulay module if $\operatorname{depth}(M)=\dim(R)$.
class ModuleCat.IsMaximalCohenMacaulay [IsLocalRing R] [Small.{v} R] (M : ModuleCat.{v} R) : Prop where
depth_eq_dim : IsLocalRing.depth M = ringKrullDim R
As we know from subsection  3.3 that $\operatorname{depth}(M)\leq\dim(M)\leq\dim(R)$, maximal Cohen–Macaulay modules can be viewed as modules over $R$ with the maximum possible $\operatorname{depth}$. In particular, a maximal Cohen–Macaulay module is Cohen–Macaulay.W
The following theorem is the main result of this subsection.W
Theorem 5.7 ([22, 00NT]).
For regular local ring $R$, any finitely generated maximal Cohen–Macaulay module over $R$ is free.
theorem free_of_isMaximalCohenMacaulay_of_isRegularLocalRing [IsRegularLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] [M.IsMaximalCohenMacaulay] : Module.Free R M
Proof.
We proceed by induction on $\dim(R)$. If $\dim(R)=0$, then $R$ is a field and $M$ is free. Assume the statement for $n$, and let $\dim(R)=n+1$. Choose $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$; by Corollary  5.2, the quotient $R/(x)$ is a regular local ring of dimension $n$. To show $x$ is $M$-regular it suffices to prove that for every $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $x\notin\mathfrak{p}$. By Corollary  3.13, $\dim(R/\mathfrak{p})\geq\operatorname{depth}(M)=\dim(R)=n+1$. If $x\in\mathfrak{p}$, then $n=\dim(R/(x))\geq\dim(R/\mathfrak{p})$, causing a contradiction. Hence $x$ is $M$-regular. From Theorem  3.6, we know that $\operatorname{depth}(M/xM)=\operatorname{depth}(M)-1=\dim(R)-1=n=\dim(R/(x))$, so $M/xM$ is a maximal Cohen–Macaulay module over $R/(x)$. By the induction hypothesis, $M/xM$ is free over $R/(x)$, the result then follows from Lemma  5.5. ∎
5.2.2. The Main Proof
The proof idea of Theorem  5.4 is standard. For a regular local ring $(R,\mathfrak{m},k)$ of dimension $d$ and any $R$-module $M$, repeated application of Lemma  3.5 shows that the $d$-th syzygy of $M$ in a free resolution is either 0 or maximal Cohen–Macaulay; in either case, it is projective. (See arguments in [5, Theorem 2.2.7] for details.) However, formalizing a distant syzygy and the depth calculations is cumbersome, so we use a trick to simplify the formalization.
For a regular local ring $R$, the Auslander–Buchsbaum formula implies that for a finitely generated $R$-module $M$ of finite projective dimension, the projective dimension does not exceed $\operatorname{depth}(R)=\dim(R)$. Since $\operatorname{depth}(k)=0$, if $k$ has finite projective dimension, then its projective dimension equals $\operatorname{depth}(R)=\dim(R)$. Thus, it suffices to show that every finitely generated module over a regular local ring has finite projective dimension. W
lemma projectiveDimension_ne_top_of_isRegularLocalRing [IsRegularLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] : projectiveDimension M Top.top
We then set up an auxiliary result that is easier to prove by induction on $i\in\mathbb{N}$, using only single-step dimension shifting.W
Proposition 5.8.
For any $i\in\mathbb{N}$, if $\operatorname{depth}(M)+i\geq\dim(R)$ then $M$ has finite projective dimension.
lemma finite_projectiveDimension_of_isRegularLocalRing_aux [IsRegularLocalRing R] [Small.{v} R]
(M : ModuleCat.{v} R) [Module.Finite R M] (i : ℕ) :
IsLocalRing.depth M + i ringKrullDim R n, HasProjectiveDimensionLE M n
Proof.
For $i=0$, $\operatorname{depth}(M)\geq\dim(R)$ implies equality, so $M$ is maximal Cohen–Macaulay and hence free by Theorem  5.7; thus $M$ has finite projective dimension. Assume the claim for $i$. For $i+1$, suppose $M\neq 0$, take a short exact sequence $0\to N\to F\to M\to 0$ with $F$ free and finitely generated. By Lemma  3.5, $\operatorname{depth}(N)+i\geq\min\{\operatorname{depth}(F),\operatorname{depth}(M)+1\}+i=\min\{\operatorname{depth}(F)+i,\operatorname{depth}(M)+i+1\}$. Since $\operatorname{depth}(F)=\operatorname{depth}(R)=\dim(R)$, the right-hand side is at least $\dim(R)$, so $\operatorname{depth}(N)+i\geq\dim(R)$. By the induction hypothesis, $N$ has finite projective dimension; hence, so does $M$. This completes the induction. ∎
The proof only involves dimension shifting of one term, making it easier to formalize. Combining all the results above yields the formal proof of Theorem  5.4.
5.3. A Weakened Version of Ferrand–Vasconcelos Theorem
We first establish a lemma about projective dimension.W
Lemma 5.9 ([5, Lemma 1.3.5]).
For a finitely generated module $M$ over a Noetherian local ring $R$ and for $x$ that is both $R$-regular and $M$-regular, we have $\operatorname{projdim}_{R}(M)=\operatorname{projdim}_{R/(x)}(M/xM)$.
lemma projectiveDimension_eq_quotient
[Small.{v} R] [IsLocalRing R] [IsNoetherianRing R] (M : ModuleCat.{v} R) [Module.Finite R M]
(x : R) (reg1 : IsSMulRegular R x) (reg2 : IsSMulRegular M x) (mem : x maximalIdeal R) :
projectiveDimension.{v} M = projectiveDimension.{v} (ModuleCat.of (R Ideal.span {x}) (QuotSMulTop x M))
Proof.
Since projective and free coincide over a local ring, Lemma 5.5 implies that $M$ is projective over $R$ if and only if $M/xM$ is projective over $R/(x)$. Using dimension shifting, it suffices to reduce the general equality of projective dimensions to the equivalence of projectivity on both sides. Concretely, take a short exact sequence $0\rightarrow N\xrightarrow{f}R^{n}\xrightarrow{g}M\rightarrow 0$. Tensoring with $R/(x)$ yields an exact sequence $N/xN\rightarrow\rightarrow(R/(x))^{n}\rightarrow M/xM\rightarrow 0$ with the obvious transition maps. Any element $y$ in the kernel of the map $N\rightarrow R^{n}\rightarrow(R/(x))^{n}$ lifts to an element of $N$ of the form $y=x\cdot z$. Since $x\cdot g(z)=g(x\cdot z)=0$, we have $z\in N$. Thus the kernel of $N\rightarrow N/xN\rightarrow(R/(x))^{n}$ equals $xN$, so the induced map $N/xN\rightarrow(R/(x))^{n}$ is injective. Because $\operatorname{projdim}_{R}(N)=\operatorname{projdim}_{R}(M)-1$ and $\operatorname{projdim}_{R/(x)}\ N/xN=\operatorname{projdim}_{R/(x)}\ M/xM-1$, and since $x$ remains $R$-regular and $N$-regular (as $N$ is submodule of a free module), the equality of projective dimensions follows. (The exactness is essentially a computation of $\operatorname{Tor}_{R}^{1}(R/(x),M)$.) ∎
Returning to the main thread, the following theorem will directly imply the desired result.
Theorem 5.10 (Ferrand–Vasconcelos, [5, Theorem 2.2.8]).
For an ideal $I$ of a Noetherian local ring $R$, if $I$ has finite projective dimension and $I/I^{2}$ is free over $R/I$, then $I$ is generated by a regular sequence.
If the residue field $k$ of a Noetherian local ring $(R,\mathfrak{m},k)$, has finite projective dimension, then so does $\mathfrak{m}$; applying the theorem to $\mathfrak{m}$ shows that $\mathfrak{m}$ is generated by a regular sequence. Since any regular sequence contained in $\mathfrak{m}$ has length at most $\operatorname{depth}(R)\leq\dim(R)$, this implies that $R$ is regular.
However, proving the full Ferrand–Vasconcelos theorem requires showing that any ideal I with a finite free resolution contains an R‑regular element, as shown in [5, Corollary 1.4.7], which in turn requires a theory of rank for free resolutions. As the theory of rank in mathlib is not yet fully developed, the complete proof is currently out of reach. The key observation is that the induction in the proof of Ferrand–Vasconcelos theorem still works if we only consider the maximal ideal. Thus, we restrict our attention to the unique maximal ideal and prove the result for that case; this already yields regularity when the residue field has finite projective dimension.W
Theorem 5.11.
For Noetherian local ring $(R,\mathfrak{m})$, if $\mathfrak{m}$ has finite projective dimension then it is generated by a regular sequence.
theorem generate_by_regular [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R]
(h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
rs : List R, IsRegular R rs Ideal.ofList rs = maximalIdeal R
Proof.
The existence of a regular element in the maximal ideal, assuming the maximal ideal has finite projective dimension and $R$ is not a field, follows from the Auslander–Buchsbaum formula (Theorem  3.8).W
Lemma 5.12.
For Noetherian local ring $(R,\mathfrak{m})$ which is not a field, $\mathfrak{m}$ contains an $R$-regular element.
Proof.
Suppose there are no regular elements in $\mathfrak{m}$. Then $\operatorname{depth}(R)=0$. If $k$ has finite projective dimension, the Auslander–Buchsbaum formula (Theorem  3.8) gives $\operatorname{projdim}(k)+\operatorname{depth}(k)=\operatorname{depth}(R)=0$. Since $\operatorname{depth}(k)=0$, we get $\operatorname{projdim}(k)=0$, so $k$ is projective hence free; thus $\mathfrak{m}=0$, contradicting the assumption that $R$ is not a field. ∎
Using prime avoidance applied to $\mathfrak{m}^{2}$ and $\mathrm{Ass}(R)$, we can further find an $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$.W
Lemma 5.13.
For Noetherian local ring $(R,\mathfrak{m})$ which is not a field, there exists an $R$-regular element in $\mathfrak{m}\setminus\mathfrak{m}^{2}$.
lemma exist_isSMulRegular_of_exist_hasProjectiveDimensionLE
[IsLocalRing R] [IsNoetherianRing R] [Small.{v} R] (nebot : maximalIdeal R ⊥)
(h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
x maximalIdeal R, x maximalIdeal R ^ 2 IsSMulRegular R x
Returning to the main proof, we proceed by induction on the span rank (minimal number of generators) of $\mathfrak{m}$. If this number is 0, $\mathfrak{m}=0$ is generated by the empty sequence. Assume the statement for $n$, and consider the case $n+1$. By Lemma  5.13, we can choose an $R$-regular element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$. Denote by $\tilde{\mathfrak{m}}$ the maximal ideal of $R/(x)$; its span rank is exactly $n$. To apply the induction hypothesis to $R/(x)$, it remains to show $\tilde{\mathfrak{m}}$ has finite projective dimension over $R/(x)$. Let $\overline{J}$ be a direct sum complement of $\mathrm{span}\{\overline{x}\}$ in $\mathfrak{m}/\mathfrak{m}^{2}$, and let $J$ be the preimage of $\overline{J}$ in $\mathfrak{m}$. Then $J+(x)=\mathfrak{m}$. For any $y\in J\cap(x)$, write $y=xz$. Because $\overline{y}\in\mathrm{span}\{\overline{x}\}\cap\overline{J}=0$ in $\mathfrak{m}/\mathfrak{m}^{2}$, we get $z\in\mathfrak{m}$ and hence $y\in x\mathfrak{m}$. Thus $J\cap(x)\subseteq x\mathfrak{m}$. From this one obtains maps whose composition is the identity:
$\tilde{\mathfrak{m}}\cong(J+(x))/(x)\cong J/(J\cap(x))\rightarrow\mathfrak{m}/x\mathfrak{m}\rightarrow\tilde{\mathfrak{m}}$
Therefore $\tilde{\mathfrak{m}}$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$ as an $R/(x)$ module. By Lemma  5.9, $\operatorname{projdim}_{R/(x)}\ \mathfrak{m}/x\mathfrak{m}=\operatorname{projdim}_{R}\mathfrak{m}<\infty$, so $\tilde{\mathfrak{m}}$ has finite projective dimension over $R/(x)$. Applying the induction hypothesis, $\tilde{\mathfrak{m}}$ is generated by an $R/(x)$-regular sequence $rs$. Prepending $x$ yields $x::rs$, an $R$-regular sequence generating $\mathfrak{m}$. This completes the induction. ∎
We conclude this subsection with the following two direct corollaries of Theorem 5.11.WW
Theorem 5.14.
For a Noetherian local ring $(R,\mathfrak{m})$, if $\mathfrak{m}$ has finite projective dimension then $R$ is regular.
theorem IsRegularLocalRing.of_maximalIdeal_hasProjectiveDimensionLE [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
IsRegularLocalRing R
Theorem 5.15.
Noetherian local ring having finite global dimension is regular.
theorem IsRegularLocalRing.of_globalDimension_lt_top [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R]
(h : globalDimension.{v} R < Top.top) : IsRegularLocalRing R
5.4. The Final Conclusions and Related Results
5.4.1. Auslander–Buchsbaum–Serre Criterion
With the results in Subsections  5.2 and  5.3, we can finally state and prove the Auslander–Buchsbaum–Serre criterion, following the original version in [1], as follows.W
Theorem 5.16.
A Noetherian local ring $R$ is regular if and only if $\operatorname{gldim}(R)$ is finite.
theorem Auslander_Buchsbaum_Serre [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R] :
IsRegularLocalRing R globalDimension.{v} R < Top.top
As noted in Lemma  2.14, the global dimension of a ring $R$ equals the supremum of the global dimensions of the localizations $R_{\mathfrak{p}}$ as $\mathfrak{p}$ ranges over the prime ideals. Hence if $R$ has finite global dimension, then each $R_{\mathfrak{p}}$ does as well. This yields the following useful corollary.W
Corollary 5.17.
For regular local ring $R$ and $\mathfrak{p}$ a prime ideal of $R$, the localization $R_{\mathfrak{p}}$ is regular.
lemma IsRegularLocalRing.of_isLocalization [IsRegularLocalRing R] (p : Ideal R) [p.IsPrime]
(S : Type*) [CommRing S] [Algebra R S] [IsLocalization.AtPrime S p] : IsRegularLocalRing S
5.4.2. Regular Ring
We also developed the basic theory of regular rings.W
Definition 5.18.
A ring $R$ is a regular ring if $R_{\mathfrak{p}}$ is a regular local ring for every prime ideal $\mathfrak{p}$.
class IsRegularRing : Prop where
localization_isRegular : p : Ideal R, (_ : p.IsPrime), IsRegularLocalRing (Localization.AtPrime p)
There is analogous statement of Theorem  5.4 for regular rings.W
Theorem 5.19.
For regular ring $R$, $\operatorname{gldim}(R)=\dim(R)$.
theorem IsRegularRing.globalDimension_eq_ringKrullDim [Small.{v} R] [IsRegularRing R] [IsNoetherianRing R] :
globalDimension.{v} R = ringKrullDim R
Proof.
This is a direct consequence of Lemma  2.14 and Theorem  5.4. ∎
Since localization of a regular local ring at a prime is regular, regularity of a ring can be verified on maximal ideals only.W
Theorem 5.20.
A ring $R$ is regular if and only if $R_{\mathfrak{m}}$ is a regular local ring for every maximal ideal $\mathfrak{m}$.
lemma isRegularRing_of_localization_maximal_isRegularLocalRing
(h : m : Ideal R, (_ : m.IsMaximal), IsRegularLocalRing (Localization.AtPrime m)) :
IsRegularRing R
Finally, we are able to prove that a regular local ring is a regular ring.W
5.4.3. Hilbert’s Syzygy Theorem
With all the previous results in this section we are able to prove a weak version of the Hilbert’s Syzygy Theorem. W
Theorem 5.21 (Hilbert’s Syzygy,[16, Corollary 18.1]).
For a field $k$, $\operatorname{gldim}(k[x_{1},\cdots,x_{n}])=n$.
theorem Hilberts_Syzygy (k : Type u) [Field k] [Small.{v} k] (n : ℕ) :
globalDimension.{v} (MvPolynomial (Fin n) k) = n
However due to the current setup, we still have some distance to the original theorem first stated in [11]. To prove the weak version above, it only remains to show that a polynomial ring over a regular ring is regular.W
Theorem 5.22.
For regular ring $R$, $R[X]$ is regular.
theorem Polynomial.isRegularRing_of_isRegularRing [IsRegularRing R] : IsRegularRing R[X]
Proof.
Because both regularity and the Cohen–Macaulay property are local conditions, the proof reduces similarly to the following local case: for a regular local ring $(R,\mathfrak{m},k)$ and a prime ideal $\mathfrak{p}$ of $R[X]$ with $\mathfrak{p}\cap R=\mathfrak{m}$, show that $R[X]_{\mathfrak{p}}$ is a regular local ring.W It suffices to show that $\mathfrak{p}$ can be generated by $0pt(\mathfrak{p})=\dim(R[X]_{\mathfrak{p}})$ elements, these generators then generate the maximal ideal of $R[X]_{\mathfrak{p}}$. If $\mathfrak{p}=\mathfrak{m}[X]$, it is generated by the minimal generators of $\mathfrak{m}$. Since $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})=\dim(R)$, the localization $R[X]_{\mathfrak{p}}$ is regular. If $\mathfrak{m}[X]\subsetneq p$, the image of $\mathfrak{p}$ in $k[X]$ is generated by a single polynomial $\overline{f}$. Lift $\overline{f}$ to $f\in R[X]$, then $\mathfrak{p}$ is generated by $f$ together with the minimal generators of $\mathfrak{m}$. As $0pt(\mathfrak{p})=0pt(\mathfrak{m})+1=\dim(R)+1$, the localization $R[X]_{\mathfrak{p}}$ is regular in this case as well. ∎
6. Discussion
6.1. Remarks About Implementation
In the formalization of category theory in Lean4, one often needs to carefully consider universe levels, for example the universe level of $\operatorname{Ext}$ groups. In our implementation, for constructions on general abelian categories such as projectiveDimension and globalDimension, we use the standard universe level or leave the HasExt instance as parameters following previous definitionsW. When working on the category of $R$-modules in universe $v$, if we assume $R$ is $v$-small (denoted [Small.{v} R]), the category has enough projectives, which allows us to obtain $\operatorname{Ext}$ in universe level vvW. This is also the only universe level of $\operatorname{Ext}$ we use throughout the development. The reason is that users should not have to specify the universe parameter when considering $\operatorname{Ext}$ types, so a given category should have at most one HasExt instance. In this setup, the universe level v is therefore the only reasonable choice.
The remaining nuisance is the universe of the modules on which we take $\operatorname{Ext}$. Since the commutation of $\operatorname{Ext}$ and Ulift is not yet fully formalized yet, some definitions still partially depend on a choice of universe, such as moduleDepth, projectiveDimension and globalDimension. Currently we only know moduleDepth is invariant of universe in some good conditions by using the regular sequence characterization. However, as we know that $\operatorname{Ext}$ and Ulift do commute (up to isomorphism), it would be easy to prove that moduleDepth and projectiveDimension is invariant under arbitrary linear isomorphisms once that fact is formalized. As for globalDimension, from Lemma  2.13 we know that it suffices to consider the supremum over finitely generated modules. Notice that when assuming Small.{v} R and Small.{v’} R, there is a correspondence between finitely generated modules in ModuleCat.{v} R and ModuleCat.{v’} R. The globalDimension would be invariant when assuming the ring is small.
In our interaction with some categorical property and its corresponding versions stated in non-categorical language, for example Module.Projective, some related results had not previously taken universe levels into sufficient consideration. Fortunately, we fixed them when proceeding with our project.
6.2. Related Works
As far as we are aware, in other theorem provers the concepts of derived categories and derived functors are often not formalized, let alone theories built on them such as $\operatorname{depth}$ and homological dimension. In Lean4, this work also represents a first step toward applications of homological methods in commutative algebra.
Before this work, formalizations of commutative algebra in Lean4 have primarily been motivated by applications in algebraic geometry and algebraic number theory. For instance, the formalization of homogeneous ideals originated from the formalization of projective space [24]; developments on Dedekind domains and class groups were motivated by their roles in algebraic number theory [4]; and it is the formalization of local fields that furthered the formalization of valuation rings [7]. In contrast, the present work undertakes a systematic formalization of commutative algebra itself in combination with homological algebra, positioning it as a foundational body of theory upon which algebraic geometry and algebraic number theory can subsequently rely.
For formalizations of commutative and homological algebra in Lean4 directly related to this project, as mentioned in Subsection  2.3, there is a project that formalized that the Krull dimension of a polynomial ring over a Noetherian ring increases by exactly one W. That proof proceeds by showing (a version of) [16, Theorem 19]W. We would like to thank Andrew Yang here for his efforts in commutative algebra scattered throughout mathlib, which greatly benefited this work. All the results about derived categories and $\operatorname{Ext}$ in mathlib used in this project originate from the Liquid Tensor Experiment development [6] and were ported to Lean4 [19]. Even though we almost only used the definition of $\operatorname{Ext}$ and its long exact sequence, the work done at the level of derived categories has already laid a solid foundation for the subsequent formalization.
6.3. Future Works
Building on the development of Cohen–Macaulay rings and regular local rings, there are several natural directions for future works.
Cohen–Macaulay Modules over Arbitrary Rings
For a non-local ring $R$, one can define Cohen–Macaulay modules as follows: an $R$-module $M$ is Cohen–Macaulay if for every prime ideal $\mathfrak{p}$ of $R$, the localized module $M_{\mathfrak{p}}$ is Cohen–Macaulay over $R_{\mathfrak{p}}$.
Miracle Flatness
For a local homomorphism $R\to S$ with $R$ a regular local ring and $S$ a Cohen–Macaulay local ring, if $\dim S=\dim R+\dim(S/\mathfrak{m}_{R}S)$, then $R\to S$ is flat. This statement is known as miracle flatness (see [22, 00R4]). After the developments in this project, most of the required techniques are already formalized or are only a small step away. The remaining obstacle is formalizing a variant of the local criterion for flatness:
Lemma 6.1 ([22, 00ML]).
Let $R\rightarrow S$ be a local homomorphism of Noetherian local rings. Let $I\neq R$ be an ideal in R. Let M be a finite $S$-module. If $\operatorname{Tor}_{R}^{1}(M,R/I)=0$ and $M/IM$ is flat over $R/I$, then $M$ is flat over $R$.
Criterion of Normality
We use the following standard notation for a ring $R$:
  • ($S_{k}$) : $\operatorname{depth}(R_{\mathfrak{p}})\geq min\{k,0pt(\mathfrak{p})\}$ for all $\mathfrak{p}\in\operatorname{Spec}(R)$
  • ($R_{k}$) : if $\mathfrak{p}\in\operatorname{Spec}(R)$ and $0pt(\mathfrak{p})\leq k$, $R_{\mathfrak{p}}$ is regular.
Given our formalization of $\operatorname{depth}$, we can state the criterion of normality as follows: for a Noetherian ring $R$, $R$ is normal if and only if it satisfies ($S_{2}$) and ($R_{1}$) (see [16, Theorem 39]). Formalizing this criterion would be an interesting project that will develop many results about normal domains along the way.
Other Versions of Hilbert’s Syzygy Theorem
In Subsection  5.4.3, we proved that the global dimension of $k[x_{1},\dots,x_{n}]$ is $n$, which is only slightly stronger than the statement that every finitely generated module over $k[x_{1},\dots,x_{n}]$ has a projective resolution of length nn—the weakest version of Hilbert’s Syzygy Theorem. These projective resolutions can indeed be upgraded to free resolutions using the Quillen–Suslin theorem (see [18][23]) or other constructions specific to polynomial rings over fields (see [8, Corollary 19.8]). Additionally, there is a graded version of this theorem (see [8, Corollary 19.7]); however, with the current setup, it may not be feasible to formalize the graded version.